### v4

# Higgs Inflation Model with Small Non-Minimal Coupling Constant

Abstract
In the proposed model of the "Higgs field$+$gravity" system, the self-consistency of the system of equations obtained from the original action has the form of an algebraic equation (constraint) defining the scalar $\zeta$ as a function of the Higgs field $\varphi$ and its first derivatives. The scalar $\zeta$ is present in all equations of motion and has a significant effect on the dynamics of the fields. After the transition in the equations of motion to the Einstein frame with the spatially flat Friedmann metric, it is convenient to describe the resulting system of equations using the action $S_{eff}$ and the Lagrangian $L_{eff}$, which we call the effective action and the effective Lagrangian and from which these equations can be obtained. Due to the constraint, the original model parameters are converted in $L_{eff}$ into $\varphi$-dependent classical effective parameters. In particular, the effective potential $U_{eff}(\varphi)$ in $L_{eff}$ has the form $U_{eff}=\frac{\lambda}{4\xi^{2}}M_{P}^{4}\cdot F(\varphi)\cdot\tanh^{4}\bigl{(}\frac{\sqrt{\xi}\varphi}{M_{P}}\bigr{)}$, where $F(\varphi)$ is a smooth function equal to $F(\varphi)\approx\frac{1}{2}$ for $\varphi\gg\sqrt{3/8}M_{P}$. The constant $\xi$ of non-minimal coupling to scalar curvature can be chosen as small as desired. If $\xi=1/6$, then to ensure agreement with CMB observational data, the Higgs field self-coupling parameter $\lambda$ in the original action must be of the order of $\sim 10^{-11}$. During cosmological evolution after the end of inflation, the decrease of $\varphi$ leads to a change in the sign of the effective Higgs mass term in $L_{eff}$. This TMSM effect provides an answer to the mystery of the Higgs potential structure and leads to spontaneous symmetry breaking. As $\varphi$ approaches VEV, the scalar function $\zeta(\varphi)$ changes in such a way that the classical effective self-coupling parameter $\lambda_{eff}(\zeta(\varphi))$ increases by 10 orders of magnitude compared to $\lambda$, which is necessary for the implementation of the GWS theory. Applying the model to the very beginning of the classical evolution of the Universe shows that under certain initial conditions, cosmological dynamics can begin with a "pathological" and even phantom regime preceding inflation. However, if evolution begins with normal dynamics, then it proceeds only as inflation, and the problem of initial conditions for the onset of inflation does not arise.
1 Introduction
Theoretical studies and accumulated observational data have led to the generally accepted view that cosmological inflation was the initial stage of the classical evolution of the Universe. Despite this, one of the main problems in modern high-energy physics and cosmology remains the question of what physical object or effect drives cosmological inflation and subsequent graceful exit from it. In the bewildering variety of existing inflation models [1], the mysterious degree of freedom responsible for inflation is ultimately described by a scalar field, usually called the inflaton. If we do not want to be satisfied with the framework of the phenomenological description, but seek to understand the nature of the inflaton based on the existing experimentally confirmed theory of particle physics, then the most suitable and, perhaps, the only candidate is the $SU(2)\times U(1)$ gauge-invariant electroweak standard model (SM). This means that the SM Higgs field must play the role of an inflaton. The corresponding scenarios are realized in the metric-Higgs inflation models [2]-[11], as well as in the alternative approach using the Palatini formulation and developed in refs. [12]-[19].
To compare all existing Higgs inflation models with the model proposed in this paper, we begin with a brief description of the main steps of their construction at the tree level. The electroweak SM at the tree level is defined in Minkowski space. To use the SM in cosmology there is a need to reformulate it in a general covariant form to allow gravity to be included in the model. The next steps are to add a nonminimal coupling of the Higgs field to the scalar curvature, followed by the transition to the Einstein frame in the original action and a redefinition of the Higgs field leading to a Lagrangian with a canonically normalized field. This combination of particle physics and gravity, which we will call SM+Gravity, is the basis for how the Higgs sector of the electroweak SM is commonly used to study cosmology at energies well above the electroweak energy scale. However, the resulting models are suitable for describing inflation in accordance with the CMB data [20], [21] only if the constant of non-minimal coupling to scalar curvature is unnaturally huge: depending on the type of model, metric or Palatini, the coupling constant varies in the range $10^{4}\lesssim\xi\lesssim 10^{8}$. Nevertheless, apart from some effects due to nonminimal coupling (see, for example, review [10]), in the SM+Gravity approach to Higgs inflation models at the tree level, the question of finding a reasonable connection between low-energy particle physics and physics at the inflation energy scale remains open. Moreover, the need to take into account quantum corrections runs into problems associated with the non-renormalizability of the resulting models, which further increases doubts about the validity of the SM+Gravity approach.
In the present paper, the Higgs inflation model is formulated in the framework of the Two Measure Theory (TMT), which radically changes the approach to combining gravity with particle physics. Due to these changes, a formulation of the SM that takes into account the structure and procedure of TMT (and which we will refer to as Two-Measure Standard Model (TMSM)) can successfully overcome the above-mentioned problems of Higgs inflation.
  • Gravity is built into TMT in such a way that in all TMT models the equations derived from the principle of least action describe a self-consistent system of gravity and matter, and this system of equations, represented in the Einstein frame, is valid at any energy available to classical field theory. One of the most important consequences of this feature of TMT is that, in the context of cosmology, the TMSM we study describes what the tree-level SM looks like at various stages of cosmological evolution, which is equivalent to a description at any allowed energies. The results of studying these aspects of TMSM will be presented in the next paper [22].
  • The radical difference between TMSM and SM+Gravity is most obvious in the way these theories describe matter in Minkowski space. In SM+Gravity, it is usually assumed that somehow the zero value of the vacuum energy is already ensured, and so the cosmological constant (CC) in the Lagrangian is set to zero "manually". This approach means that models like Higgs+Gravity are formulated within a theory that is not self-sufficient in such a fundamental question. In TMSM, the description starts with a general coordinate-invariant initial action, including the gravitational terms, the vacuum-like terms and the SM matter. The description of matter in Minkowski space is only possible as a result of the limiting transition in all equations of motion to the vacuum state. The latter is possible only by choosing the appropriate integration constant ${\mathcal{M}}$, i.e. by fine-tuning. Moreover, the CC problem and the method of ensuring its required value (not necessarily zero) also turns out to be built into the TMSM in a fundamentally different way compared to the SM + Gravity: if we change the integration constant ${\mathcal{M}}$ so that instead of zero the CC has any other value, this will lead to corresponding changes not only in the Higgs field potential (which takes place in SM+Gravity), but also in the values of all TMSM parameters that are present in the equations of all SM matter fields. Therefore, despite the fact that the described fine-tuning cannot be considered a solution to the CC problem, the implementation of this fine-tuning has a significant effect on all elements of the model.
  • The main result of the present paper is the demonstration that the Higgs+gravity sector of TMSM with an arbitrarily small constant $\xi$ of non-minimal coupling to the scalar curvature allows one to construct a model of Higgs inflation that agrees well with observational CMB data. At the same time, despite the seemingly obvious expectations, there is no contradiction between this result and the fact that the TMSM also gives the correct values of the SM physical quantities at accelerator energies. Moreover, the TMSM provides a natural explanation for the mystery of the appearance of a mass term with the "wrong" sign in the Higgs potential as an effect arising during cosmological evolution due to a special property of the TMSM associated with the presence of two volume measures in the original action.
  • The second, but no less important result of this paper concerns the problem of initial conditions for the onset of inflation and the possible stages of evolution preceding inflation.
Conditions that the initial kinetic and gradient energy densities of the canonically normalised scalar field should not exceed the initial value of the potential energy density
$\rho_{kin}^{(in)}=\frac{1}{2}\dot{\varphi}_{in}^{2}\lesssim V(\varphi_{in})\quad\text{and}\quad\rho_{grad}^{(in)}=\frac{1}{2}|(\partial^{k}\varphi)_{in}(\partial_{k}\varphi)_{in}|\lesssim V(\varphi_{in})$ (1.1)
are well known as the constraints neeeded for the onset of inflation. The initial understanding of this issue was developed in the first models of chaotic inflation [23], [24]. However, later, the need to implement flat potential models [25], [26], [27], [28], [29], [2], in which inflation predictions agree well with the SMB observational data, turned the question of initial conditions into a problem. The seriousness and urgency of this problem became clear from the discussion in papers [30]-[34] . A completely new approach to the problem of initial conditions for inflation based on the TMT was proposed in ref. [35], where it was shown that the solution to the problem is closely related to the need to answer the question of what happened before inflation. Within the framework of the TMT model of ref. [35], where the inflaton field was considered independent of the known matter fields, the problem of initial conditions for inflation turned out to be related to the GBV theorem [36]-[39]. In the present paper, in which the inflaton field is identified with the Higgs field, we find that not only are the initial conditions (1.1) for inflation guaranteed to be satisfied, but an earlier, rather exotic stage of evolution preceding inflation is also possible.
From the most general point of view, comparing the SM+Gravity models with TMSM, we can say that SM+Gravity belongs to the “conventional” type of models, in which the original action contains a single volume element $dV_{g}=\sqrt{-g}d^{4}x$. In models of the Two-Measure Theory, in the integral of the primordial action, along with terms with the standard volume measure $dV_{g}=\sqrt{-g}d^{4}x$, there are terms with an alternative, metric independent volume measure $dV_{\Upsilon}$ constructed as the following 4-form using 4 scalar functions $\varphi_{a}$ ($a=1,..,4$)
$dV_{\Upsilon}=\Upsilon d^{4}x\equiv\varepsilon^{\mu\nu\gamma\beta}\varepsilon_{abcd}\partial_{\mu}\varphi_{a}\partial_{\nu}\varphi_{b}\partial_{\gamma}\varphi_{c}\partial_{\beta}\varphi_{d}d^{4}x=4!d\varphi_{1}\wedge d\varphi_{2}\wedge d\varphi_{3}\wedge d\varphi_{4}.$ (1.2)
Here $\Upsilon$ is a scalar density, that is under general coordinate transformations with positive Jacobian it has the same transformation law as $\sqrt{-g}$. In contrast to the density $\sqrt{-g}$ of the standard volume measure $dV_{g}$, the density $\Upsilon$ of the 4-form $dV_{\Upsilon}$ is sign-indefinite and, due to its continuity, can also take zero values in the general case. A specific dynamic feature of TMT is that as a result of varying the initial action, the ratio of volume measures
$\zeta(x)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{dV_{\Upsilon}}{dV_{g}}\equiv\frac{\Upsilon}{\sqrt{-g}}$ (1.3)
appears in all equations of motion and it turns out to be $\phi$-dependent. The scalar function $\zeta$ has a key role in obtaining all the new results presented in this paper.
The organization of the paper is as follows. In section 2, after defining the primordial action [^2] In ”conventional” alternative gravity theories, the original action differs from the Einstein-Hilbert action only in the form of the Lagrangian, for example, by the presence of a non-minimal coupling with curvature. The transformation to the Einstein frame, which in conventional theories is performed in the original action, simply changes the set of variables used to describe the theory, but does not change the theory itself. In TMT, the main difference is that the volume measure contains $\Upsilon$, although a modification of the Lagrangian is also possible. Therefore, the original TMT action differs from the Einstein-Hilbert action in the form of the Lagrangian density, and to bring the TMT action to the Einstein frame, the conformal transformation must exclude from the gravitational term of the action not only the non-minimal coupling, but also $\Upsilon$. In a theory obtained in this way, the constraint determining $\zeta$ cannot arise, that is, by acting in this way we would be dealing with a different theory. It is therefore incorrect to use the term ”Jordan frame” in the usual sense for the set of variables in the original action of TMT. To emphasize this distinction, we will use the terms ”primordial variables”, ”primordial model parameters”, ”primordial Lagrangian” and ”primordial action” instead of ”variables in the original frame”, ”original model parameters”, ”original Lagrangian” and ”original action”. and justifying the need to add a new type of vacuum-like term related to the volume measure density $\Upsilon$, we perform all the steps required by the TMT procedure. Having found the equations and represented them in the Einstein frame in $\zeta$-dependent form, we, using the constraint, arrive at the TMT-effective equations, potential, Lagrangian and action [^3] Note that this use of the term ”effective” refers to the classical description and, for example, is in no way connected with the usual use of the term ”effective potential” calculated taking into account quantum corrections. for the Higgs sector of the TMSM (where all gauge field and fermion contributions are omitted). Their analysis shows that the obtained (classical) TMT-effective action is applicable in the entire admissible energy spectrum: from the energies of the electroweak SM to the scales of inflation. In section 3, after choosing the model parameters that provide the required flatness of the TMT effective potential for values of the canonically normalized scalar field $\varphi>6M_{P}$, we study a simplified model showing its agreement with the observational CMB data. The need to consider a simplified model (which will be justified in section 6) is due to the fact that the found TMT-effective action has a structure reminiscent of the K-essence models commonly used to describe the late time cosmic acceleration [40]-[43]. Important details of the shape of the TMT-effective potential and the behavior of the function $\zeta(\varphi)$ during the entire cosmological evolution are analyzed in section 4. This allows us to understand the role of cosmological evolution in the appearance of the negative mass term in the Higgs potential. Section 5, which is partly related to the paper [35], focuses on the effects dictated by the need to ensure that cosmological solutions (and their initial conditions) must keep the $\zeta$ function positive. In section 6 we study the influence of the K-essence type structure on the earliest stages of classical evolution of the Universe. We show that the K-essence type structure may be responsible for the existence of the earliest stages of evolution, which have "pathological" dynamics. However, if evolution starts with "normal" dynamics, then it proceeds only as inflation, and the problem of initial conditions for the onset of inflation does not arise. Finally, it is shown that the K-essence type structure does not change the results of the simplified version of the model obtained in section 3, as far as the slow-roll inflation stage is concerned. However, the K-essence type structure may have some observable effect in the late stage of inflation. In section 7 we focus on discussing two fundamentally important features of TMSM that provide the main results of the paper.
2 Higgs sector of the Standard Model in Two-Measure Theory.
General view in the context of cosmology
2.1 Action
In this paper we will study the TMSM with the following primordial action
$S_{primordial}=\int d^{4}x(\sqrt{-g}+\Upsilon)\left(-\frac{M_{P}^{2}}{2}\right)\Biggl{(}1+\xi\frac{2|H|^{2}}{M_{P}^{2}}\Biggr{)}R(\Gamma,g)+S_{H}+S_{vac}$ (2.1)
Here $M_{P}$ is the reduced Planck mass. We will proceed in Palatini’s formulation, where $\Gamma$ denotes the affine connection, and $R(\Gamma,g)=g^{\mu\nu}R_{\mu\nu}(\Gamma)$, $R_{\mu\nu}(\Gamma)=R^{\lambda}_{\mu\nu\lambda}(\Gamma)$ and $R^{\lambda}_{\mu\nu\sigma}(\Gamma)\equiv\Gamma^{\lambda}_{\mu\nu,\sigma}+\Gamma^{\lambda}_{\gamma\sigma}\Gamma^{\gamma}_{\mu\nu}-(\nu\leftrightarrow\sigma)$. $H$ is the Higgs isodoublet. Following what was said in the Introduction, all contributions from gauge fields and fermions are omitted. In our notations, in a theory with only the volume element $\sqrt{-g}d^{4}x$, the model of massless scalar field non-minimally coupled to scalar curvature would be conformally invariant if the parameter $\xi$ were equal to $\xi=-\frac{1}{6}$. The primordial action for the Higgs field $H$ is chosen as follows
$S_{H}=\int d^{4}x\left[(b_{k}\sqrt{-g}-\Upsilon)g^{\alpha\beta}\left(\partial_{\alpha}H\right)^{{\dagger}}\partial_{\beta}H-(b_{p}\sqrt{-g}+\Upsilon)\lambda|H|^{4}-(b_{p}\sqrt{-g}-\Upsilon)m^{2}|H|^{2}\right],$ (2.2)
The model parameters $b_{k}>0$, $b_{p}>0$, $\lambda>0$ and $m^{2}>0$ are chosen to be positive. Given that the FLRW Universe is a model resulting from cosmological averaging, we can consider the corresponding average of $H$. Let $\phi/\sqrt{2}$ denote the only non-zero component of $H$ obtained by cosmological averaging, and we assume that $\phi$ plays the role of the inflaton.
Contribution of the vacuum-like terms to the primordial action is defined by
$S_{vac}=\int d^{4}x\left(-\sqrt{-g}V_{1}-\frac{\Upsilon^{2}}{\sqrt{-g}}V_{2}\right),$ (2.3)
If the first term in eq.(2.3) was present in Einstein’s GR, $V_{1}$ would be a cosmological constant. The term with $V_{2}$ was first introduced in Ref. [46]. The first reason for adding the term with $V_{2}$ is that the term $\propto\Upsilon=\zeta(x)\sqrt{-g}$, which can be expected as the contribution of quantum-gravitational effects to a vacuum-like action, does not contribute to the equations of motion. Therefore, the next in powers of $\zeta$ vacuum-like term can be of the form $\propto\zeta^{2}\sqrt{-g}=\Upsilon^{2}/\sqrt{-g}$. There are also more pragmatic reasons. It turns out that thanks to this term, the corresponding TMT-effective potential and the TMT-effective action acquire fundamentally new important properties without which the goals of this paper could not be achieved.
2.2 Equations of motion in the Einstein frame
Following the prescription of the TMT procedure we start with the equations of motion minimizing the primordial action (2.1). Varying the action with respect to scalar functions $\varphi_{a}$ of which $\Upsilon$ is built we get
$B^{\mu}_{a}\partial_{\mu}\left[-\frac{M_{P}^{2}}{2}\Omega R(\Gamma,g)-\frac{1}{2}g^{\alpha\beta}\phi_{,\alpha}\phi_{,\beta}-\frac{\lambda}{4}\phi^{4}+\frac{1}{2}m^{2}\phi^{2}-2\zeta V_{2}\right]=0,$ (2.4)
$\Omega=\Omega(\phi^{2})=1+\xi\frac{\phi^{2}}{M_{P}^{2}},$ (2.5)
$B^{\mu}_{a}=\varepsilon^{\mu\nu\alpha\beta}\varepsilon_{abcd}\partial_{\nu}\varphi_{b}\partial_{\alpha}\varphi_{c}\partial_{\beta}\varphi_{d},$ (2.6)
and the scalar $\zeta(x)$ defined by eq.(1.3) appears. Since $Det(B^{\mu}_{a})\propto\Upsilon^{3}$ it follows that if
$\quad\textsf{everywhere}\quad\Upsilon(x)\neq 0,$ (2.7)
the equality
$-\frac{M_{P}^{2}}{2}\Omega R(\Gamma,g)-\frac{1}{2}g^{\alpha\beta}\phi_{,\alpha}\phi_{,\beta}-\frac{\lambda}{4}\phi^{4}+\frac{m^{2}}{2}\phi^{2}-2\zeta V_{2}={\mathcal{M}}$ (2.8)
must be satisfied, where ${\mathcal{M}}$ is a constant of integration with the dimension of $(mass)^{4}$.
Variation with respect to $g^{\mu\nu}$ yields the equation
$(1-\zeta)\left(-\frac{M_{P}^{2}}{2}\right)\Omega R_{\mu\nu}(\Gamma)+(b_{k}-\zeta)\frac{1}{2}\phi_{,\mu}\phi_{,\nu}$
$-\frac{1}{2}g_{\mu\nu}\left[-\frac{M_{P}^{2}}{2}\Omega R(\Gamma,g)+\frac{b_{k}}{2}g^{\alpha\beta}\phi_{,\alpha}\phi_{,\beta}-b_{p}\frac{\lambda}{4}\phi^{4}-b_{p}\frac{m^{2}}{2}\phi^{2}-V_{1}+\zeta^{2}V_{2}\right]=0$ (2.9)
Using the trace of eq.(2.9) one can eliminate the term $-\frac{1}{2}M_{P}^{2}\Omega R(\Gamma,g)$ from eq.(2.8). As a result, we find that for these equations to be consistent, it is necessary that the scalar function $\zeta(x)$ satisfy the following relation
$\zeta\left[{\mathcal{M}}-2V_{2}+\frac{\lambda}{4}\phi^{4}-\frac{m^{2}}{2}\phi^{2}\right]$ (2.10)
-- ${\mathcal{M}}+2V_{1}+(2b_{p}-1)\frac{\lambda}{4}\phi^{4}+(2b_{p}+1)\frac{1}{2}m^{2}\phi^{2}-(1+b_{k})\frac{1}{2}g^{\alpha\beta}\phi_{,\alpha}\phi_{,\beta}=0.$
It is interesting to note that during these algebraic manipulations the term $2\zeta^{2}V_{2}$ in (2.10) cancels out. Following the terminology of earlier TMT models [43]-[53], we will call this a constraint. It should be noted that in the Palatini formulation $\zeta(x)$ is not a physical degree of freedom. Therefore, when we call Eq.(2.10) a constraint, we must keep in mind that it is different in meaning from the usual constraint in the field theory models, where it describes the relationship between dynamical degrees of freedom.
Variation of the affine connection yields the equations solution of which has the following form
$\Gamma^{\lambda}_{\mu\nu}=\{^{\lambda}_{\mu\nu}\}+(\delta^{\lambda}_{\mu}\chi,_{\nu}+\delta^{\lambda}_{\nu}\chi,_{\mu}-\chi,_{\beta}g_{\mu\nu}g^{\lambda\beta}),$ (2.11)
where $\{^{\lambda}_{\mu\nu}\}$ are the Christoffel’s connection coefficients of the metric $g_{\mu\nu}$ and
$\chi(x)=\frac{1}{2}\ln\left[\bigl{(}1+\zeta\bigr{)}\Omega\right].$ (2.12)
If $\chi(x)\neq const.$ the metricity condition does not hold and consequently geometry of the space-time with the metric $g_{\mu\nu}$ is generically non-Riemannian. In this paper, we will totally ignore a possibility to incorporate the torsion tensor, which could be an additional source for the space-time to be different from Riemannian. It is easy to see that the transformation of the metric
$\tilde{g}_{\mu\nu}=\bigl{(}1+\zeta\bigr{)}\Omega g_{\mu\nu}$ (2.13)
turns the affine connection $\Gamma^{\lambda}_{\mu\nu}$ into the Christoffel connection coefficients of the metric $\tilde{g}_{\mu\nu}$ and the space-time turns into (pseudo) Riemannian.
Gravitational equations (2.9) expressed in terms of the metric $\tilde{g}_{\mu\nu}$ take the canonical GR form
$R_{\mu\nu}(\tilde{g})-\frac{1}{2}\tilde{g}_{\mu\nu}R(\tilde{g})=\frac{1}{M_{P}^{2}}T_{\mu\nu}^{(eff)}$ (2.14)
with the same Newton constant as in the original frame. Here $R_{\mu\nu}(\tilde{g})$ and $R(\tilde{g})$ are the Ricci tensor and the scalar curvature of the metric $\tilde{g}_{\mu\nu}$, respectively. Therefore, gravity becomes canonical, and the set of dynamical variables using the metric $\tilde{g}_{\mu\nu}$ can be called the Einstein frame. $T_{\mu\nu}^{(eff)}$ on the right side of the Einstein equations (2.14) is the TMT-effective energy-momentum tensor
$T_{\mu\nu}^{(eff)}=\frac{1}{\Omega}\Bigl{(}\frac{b_{k}-\zeta}{1+\zeta}\phi_{,\mu}\phi_{,\nu}+\tilde{g}_{\mu\nu}X_{\phi}\Bigr{)}+\tilde{g}_{\mu\nu}U_{eff}^{(tree)}(\phi,\zeta(\phi);{\mathcal{M}}),$ (2.15)
where $X_{\phi}=\frac{1}{2}\tilde{g}^{\alpha\beta}\phi_{,\alpha}\phi_{,\beta}$ and
$U_{eff}^{(tree)}(\phi,\zeta;{\mathcal{M}})=\frac{1}{\Omega^{2}}\left[\frac{{\mathcal{M}}-V_{1}-V_{2}+(1-b_{p})\frac{\lambda}{4}\phi^{4}-(1+b_{p})\frac{m^{2}}{2}\phi^{2}}{(1+\zeta)^{2}}+V_{2}\right]$ (2.16)
is the $\zeta$-dependent form of the TMT-effective potential [^4] The terms “TMT-effective energy-momentum tensor” and “TMT-effective potential” are used to denote the energy-momentum tensor and potential that appear in the equations of motion after performing all the steps of the TMT procedure described above, which begins with varying the primordial TMT action and ends with the transition to the Einstein frame in the equations of motion. The superscript (tree) in $U_{eff}^{(tree)}$ is added to avoid confusion with the effective potential obtained by taking into account quantum corrections.. It is important to note once again that in Palatini’s formalism $\zeta$ is not a dynamical variable and its value is determined by the constraint (2.10), which is conveniently represented in the Einstein frame as
$\zeta=\zeta(\phi,X_{\phi};\mathcal{M})=\frac{\mathcal{M}-2V_{1}-(2b_{p}-1)\frac{\lambda}{4}\phi^{4}-(2b_{p}+1)\frac{m^{2}}{2}\phi^{2}+(1+b_{k})\Omega X_{\phi}}{\mathcal{M}-2V_{2}+\frac{\lambda}{4}\phi^{4}-\frac{m^{2}}{2}\phi^{2}+(1+b_{k})\Omega X_{\phi}}.$ (2.17)
From condition (2.7) and definition (1.3) it follows that during the entire evolution of the Universe $\zeta(x)$ cannot take a zero value and, therefore, there can only be $\zeta(x)>0$ or only $\zeta(x)<0$. Later on, when we use the constraint to study the dynamics of the $\phi$ field, it is fundamentally important to keep in mind that the constraint is a condition for the self-consistency of equations obtained only by variations of $\varphi_{a}$ and $g_{\mu\nu}$. Therefore, the origin of the restrictions imposed by the constraint on the dynamics of the $\phi$ field do not depend in any way on the dynamics of the $\phi$ field.
To see another significant difference between similar models in TMT and in the conventional theory, it is worth writing down the result of varying the scalar field $\phi$ in the primordial action (2.1):
$\frac{1}{\sqrt{-g}}\partial_{\mu}\left((b_{k}-\zeta)\sqrt{-g}g^{\mu\nu}\partial_{\nu}\phi\right)+(1+\zeta)\xi R(\Gamma,g)\phi+(b_{p}+\zeta)\lambda\phi^{3}+(b_{p}-\zeta)m^{2}\phi=0.$ (2.18)
In the Higgs inflation models in the conventional theory (both in the metric and in the Palatini formulation), the transition to the Einstein frame performed directly in the original action completely eliminates the non-minimal coupling of the scalar field to the scalar curvature. In the Two-Measure Standard Model under study, as in all TMT models, the transition to the Einstein frame (2.13) is carried out in the equations of motion and we notice that in the $\phi$-field equation (2.18) presented in the Einstein frame using (2.13), the term of the non-minimal coupling $\propto\xi R(\tilde{g})\phi$ is preserved. After passing to the Einstein frame in eq.(2.18) using (2.13), we can insert the expression for the scalar curvature $R(\tilde{g})$ obtained from the Einstein equations (2.14). Then, the $\phi$-field equation in the Einstein frame takes the following form
$\frac{1}{\sqrt{-\tilde{g}}}\partial_{\mu}\Bigg{(}\frac{b_{k}-\zeta}{(1+\zeta)\Omega}\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\partial_{\nu}\phi\Bigg{)}+\frac{b_{p}+\zeta}{(1+\zeta)^{2}\Omega^{2}}\lambda\phi^{3}+\frac{b_{p}-\zeta}{(1+\zeta)^{2}\Omega^{2}}m^{2}\phi$
$+\frac{2\xi\phi}{(1+\zeta)M_{P}^{2}\Omega^{3}}\left[-\mathcal{M}-2\zeta V_{2}-\frac{\lambda}{4}\phi^{4}+\frac{m^{2}}{2}\phi^{2}-(1+\zeta)\Omega X_{\phi}\right]=0.$ (2.19)
The procedure performed means that we are dealing with an explicitly formulated self-consistent system of the Higgs field and gravity.
2.3 Vacuum of the Higgs field in the context of cosmology
The dependence of $U_{eff}^{(tree)}(\phi,\zeta;{\mathcal{M}})$, eq.(2.16), on $\zeta$, which, in turn, changes during the entire cosmological evolution (from inflation to transition to vacuum), plays a decisive role in all the physical results of the theory. Below, the TMT-effective potential will be obtained using eqs.(2.16) and (2.17). From now on we will asume that $V_{1}<0$, $V_{2}<0$ and $(electroweak\,scale)^{4}\lll|V_{1}|\sim|V_{2}|<M_{P}^{4}$ and we will see that by choosing the integration constant $\mathcal{M}$ one can ensure that the TMT-effective potential has a minimum at some value $\phi|_{min}=\sigma\lll M_{P}$. One can treat $\sigma$ as a vacuum expectation value (VEV) of $\phi$. The corresponding value $\zeta_{v}$ in the vacuum is obtained by making use of the constraint (2.17)
$\zeta_{v}=\frac{\mathcal{M}+2|V_{1}|-(2b_{p}-1)\frac{\lambda}{4}\sigma^{4}-(2b_{p}+1)\frac{m^{2}}{2}\sigma^{2}}{\mathcal{M}+2|V_{2}|+\frac{\lambda}{4}\sigma^{4}-\frac{m^{2}}{2}\sigma^{2}}$ (2.20)
Then eq.(2.19) reduces to an equation with a canonically normalised scalar field
$\frac{1}{\sqrt{-\tilde{g}}}\partial_{\mu}\bigl{(}\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\partial_{\nu}\phi\bigr{)}+V^{\prime}_{eff}(\phi)|_{near\,vac}=0,$ (2.21)
where the derivative $V^{\prime}_{eff}(\phi)|_{near\,vac}$ in the vicinty of the vacuum has the following form
$V^{\prime}_{eff}(\phi)|_{near\,vac}$ $=$ $\frac{1}{(1+\zeta_{v})(b_{k}-\zeta_{v})}\bigl{[}(b_{p}+\zeta_{v})\lambda\phi^{2}+(b_{p}-\zeta_{v})m^{2}\bigr{]}\phi$ (2.22)
$+$ $\frac{2\xi}{(b_{k}-\zeta_{v})M_{P}^{2}}\Bigl{[}-\mathcal{M}+2\zeta_{v}|V_{2}|-\frac{\lambda}{4}\sigma^{4}+\frac{m^{2}}{2}\sigma^{2}\Bigr{]}\phi.$
Here we neglected $\frac{\xi\sigma^{2}}{M_{P}^{2}}\lll 1$ in $\Omega$. One should pay special attention that in order to obtain equation (2.21) in canonical form, it was necessary to devide eq.(2.19), considered near the vacuum, by a constant $\frac{b_{k}-\zeta_{v}}{1+\zeta_{v}}$. In secs.2.4 and 6.1 we will return to the discussion of the source of this question in more general case and will see that it is directly related to the K-essence type structure of the theory, which is responsible for the important results of the paper.
The value of the TMT effective potential in the vacuum, obtained using eqs.(2.16) and (2.20), depends on the integration constant ${\mathcal{M}}$
$U_{eff}^{(tree)}|_{(vac)}=\frac{{\mathcal{M}}^{2}-4V_{1}V_{2}+2(2b_{p}|V_{2}|+{\mathcal{M}})\frac{\lambda}{4}\sigma^{4}+2({\mathcal{M}}-2b_{p}|V_{2}|)\frac{m^{2}}{2}\sigma^{2}+\mathcal{O}(\lambda^{2}\sigma^{8})}{4\left[{|V_{1}|+|V_{2}|+\mathcal{M}}+\mathcal{O}(\lambda\sigma^{4})\right]}=\Lambda({\mathcal{M}}),$ (2.23)
where we introduced the notation $\Lambda({\mathcal{M}})$ to emphasize that with an appropriate choice of ${\mathcal{M}}$, models with a very wide range of values of the cosmological constant can be realized. The natural constraint on $\Lambda$ is $0\leq\Lambda\lesssim(electroweak\,scale)^{4}$ and it is achieved if ${\mathcal{M}}^{2}\approx 4V_{1}V_{2}$. With the choice of the integration constant [^5] By an appropriate choice of $\delta$ it is possible to ensure that $\Lambda=0$. However, this is nothing more than fine tuning, that is, in this way we do not solve the CC problem.
${\mathcal{M}}=2\sqrt{V_{1}V_{2}}(1+\delta),\qquad{\text{w}here}\qquad\delta\sim\frac{\lambda\sigma^{4}}{|V_{2}|}\lll 1,$ (2.24)
which will be used in what follows, eq. (2.20) gives
$\zeta_{v}=\sqrt{\frac{V_{1}}{V_{2}}}+\mathcal{O}\left(\frac{\lambda\sigma^{4}}{|V_{2}|}\right).$ (2.25)
We then find that near vacuum the second term in eq.(2.22) is negligible compared to the first term:
$\Bigg{|}\frac{2\xi}{(b_{k}-\zeta_{v})M_{P}^{2}}\Bigl{[}-\mathcal{M}+2\zeta_{v}|V_{2}|-\frac{\lambda}{4}\sigma^{4}+\frac{m^{2}}{2}\sigma^{2}\Bigr{]}\Bigg{|}\phi$
$=$ $\frac{2\xi}{b_{k}-\zeta_{v}}\frac{\sigma^{2}}{M_{P}^{2}}\Bigg{|}-\frac{\lambda}{4}\sigma^{2}+\frac{m^{2}}{2}\Bigg{|}\phi\lll\frac{(b_{p}+\zeta_{v})\lambda\phi^{2}+(b_{p}-\zeta_{v})m^{2}}{(1+\zeta_{v})(b_{k}-\zeta_{v})}\phi.$
In this estimate, we took into account that the implementation of Higgs inflation studied in the next section admits a large region in the parameter space that does not contradict this estimate. Then from eq.(2.22) it follows that if $\zeta_{v}>b_{p}$, spontaneous symmetry breaking (SSB) occurs with the following expression for the VEV of the classical scalar field $\phi$
$\sigma^{2}=\frac{\zeta_{v}-b_{p}}{b_{p}+\zeta_{v}}\cdot\frac{m^{2}}{\lambda}\left[1+\mathcal{O}\left(\xi\frac{\sigma^{2}}{M_{P}^{2}}\right)\right].$ (2.28)
Eqs. (2.21) and (2.22) near vacuum obviously reduce to the form of the SM equation of the Higgs field
$\frac{1}{\sqrt{-\tilde{g}}}\partial_{\mu}\bigl{(}\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\partial_{\nu}\phi\bigr{)}+\lambda_{\sigma}\phi^{3}-m_{\sigma}^{2}\phi=0$ (2.29)
and the notations [^6] We use the subscript $\sigma$ here to explicitly indicate that we are dealing with vacuum values of the TMT-effective parameters. Indeed, one might expect that, as follows from (2.19) and the constraint (2.17), $\phi$-dependent TMT-effective parameters $\lambda_{eff}$ and $m^{2}_{eff}$ appear as $\phi$ moves away from the vacuum. However, this is only partly true, namely for $\sigma<\phi\lll M_{P}$. As we will see in sections 2.4 and 4, substituting $\zeta$ from the constraint (2.17) into eq.(2.16) makes the TMT-effective potential essentially non-polynomial for $\phi\gtrsim 10^{-4}M_{P}$, which does not allow us to use the terms TMT-effective parameters $\lambda_{eff}$ and $m^{2}_{eff}$ in the general case.
$\lambda_{\sigma}=\frac{b_{p}+\zeta_{v}}{(1+\zeta_{v})(b_{k}-\zeta_{v})}\lambda,\qquad m_{\sigma}^{2}=\frac{\zeta_{v}-b_{p}}{(1+\zeta_{v})(b_{k}-\zeta_{v})}m^{2}$ (2.30)
are introduced for the values of quartic selfinteraction and squared mass parameters near the vacuum $\sigma$. Note that up to corrections of the order $\sim\mathcal{O}\bigl{(}\xi\frac{\sigma^{2}}{M_{P}^{2}}\bigr{)}$
$\sigma^{2}=\frac{\zeta_{v}-b_{p}}{b_{p}+\zeta_{v}}\cdot\frac{m^{2}}{\lambda}=\frac{m_{\sigma}^{2}}{\lambda_{\sigma}}.$ (2.31)
If we restrict ourselves to studying the TMSM near vacuum, then the parameter $m$ will remain undefined, as well as the mass parameter in the GWS theory. The presence in the relations (2.30) of factors depending on the model parameters $\zeta_{v}=\sqrt{\frac{V_{1}}{V_{2}}}$, $b_{k}$ and $b_{p}$ means that there is a very wide range of these parameters in which the required VEV SM can be ensured [^7] However, in the model studied in this paper, we completely excluded gauge bosons and fermions and, accordingly, completely ignored the need to ensure gauge invariance. A suitably extended model that modifies the above results and allows adequate comparison with the GWS theory is studied in ref. [22], where it is shown that the VEV of the Higgs field, which should be compared with the VEV in the GWS theory, is not the value of $\sigma$ given by eq.(2.31), but the value of $\sqrt{\frac{b_{k}-\zeta_{v}}{1+\zeta_{v}}}\sigma$. As will be shown in ref.[22], demonstrating this requires taking into account the need to distinguish between the description of SM in the so-called cosmological frame (CF) and in the local particle physics frame (LPPF). From this point of view, in this paper we restrict ourselves to the work in the CF.. The fundamentally new effect of the model proposed in this paper is that, thanks to these three model parameters, there is also freedom in the choice of the parameter $\lambda$. As we will see in section 3, this allows the implementation of the Higgs inflation model with a small constant $\xi$ of the non-minimal Higgs coupling to the scalar curvature.
2.4 General view of the TMSM
from the electroweak energy scale to the inflation scale
In section 2.2, TMT-effective equations in the Einstein frame were obtained. To understand the main features of TMSM at energies above the electroweak scale, it is necessary to return to these equations, taking into account the results of section 2.3 relating to what happens near the Higgs vacuum in the CF (see footnote 5). First, from the requirement that the TMT-effective cosmological constant be negligibly small, we obtained that, neglecting insignificant corrections, the integration constant is $\mathcal{M}=2\sqrt{V_{1}V_{2}}$, eq.(2.24). Secondly, the scalar $\zeta$ in vacuum is equal to $\zeta_{v}=\sqrt{\frac{V_{1}}{V_{2}}}$, eq.(2.25). Therefore the term $\mathcal{M}+2\zeta V_{2}$ in $\phi$-equation (2.19) must be replaced by $2|V_{2}|(\zeta_{v}-\zeta)$.
$\frac{1}{\sqrt{-\tilde{g}}}\partial_{\mu}\Bigg{(}\frac{b_{k}-\zeta}{(1+\zeta)\Omega}\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\partial_{\nu}\phi\Bigg{)}+\frac{b_{p}+\zeta}{(1+\zeta)^{2}\Omega^{2}}\lambda\phi^{3}+\frac{b_{p}-\zeta}{(1+\zeta)^{2}\Omega^{2}}m^{2}\phi$
$+\frac{2\xi\phi}{(1+\zeta)M_{P}^{2}\Omega^{3}}\left[-2|V_{2}|(\zeta_{v}-\zeta)-\frac{\lambda}{4}\phi^{4}+\frac{m^{2}}{2}\phi^{2}-(1+\zeta)\Omega X_{\phi}\right]=0.$ (2.32)
Note that in the vacuum the $2|V_{2}|(\zeta_{v}-\zeta)$ term is zero.
From now on, we will use the following parametrization for $V_{2}$
$|V_{2}|=(qM_{P})^{4},\quad q\leq 1.$ (2.33)
After substituting $\zeta(\phi,X_{\phi})$, eq.(2.17), into the $\zeta$-dependent form of $T_{\mu\nu}^{(eff)}$ in eqs.(2.15) and (2.16) with subsequent simple but rather lengthy algebraic calculations, the expression for the TMT-effective energy-momentum tensor reduces to
$T_{\mu\nu}^{(eff)}=\frac{K_{1}(\phi)}{\Omega}\cdot\big{(}\phi_{,\mu}\phi_{,\nu}-\tilde{g}_{\mu\nu}X_{\phi}\big{)}-K_{2}(\phi)\cdot\frac{X_{\phi}}{M_{P}^{4}}\Big{(}\phi_{,\mu}\phi_{,\nu}-\frac{1}{2}\tilde{g}_{\mu\nu}X_{\phi}\Big{)}+\tilde{g}_{\mu\nu}U_{eff}^{(tree)}(\phi),$ (2.34)
where the TMT effective potential $U_{eff}^{(tree)}(\phi)$ and functions $K_{1}(\phi)$ and $K_{2}(\phi)$ are defined by the following expressions:
$U_{eff}^{(tree)}(\phi)=\frac{q^{4}\Bigl{[}(\zeta_{v}+b_{p})\lambda\phi^{4}-2(\zeta_{v}-b_{p})m^{2}\phi^{2}\Bigr{]}+\frac{1}{M_{P}^{4}}\bigl{(}\frac{\lambda}{4}\phi^{4}-\frac{m^{2}}{2}\phi^{2}\bigr{)}^{2}}{4\Omega^{2}\left[q^{4}(1+\zeta_{v})^{2}+(1-b_{p})\frac{\lambda}{4}\frac{\phi^{4}}{M_{P}^{4}}-(1+b_{p})\frac{m^{2}\phi^{2}}{2M_{P}^{4}}\right]},$ (2.35)
$K_{1}(\phi)=\frac{2q^{4}(1+\zeta_{v})(b_{k}-\zeta_{v})+(2b_{p}+b_{k}-1)\frac{\lambda}{4}\frac{\phi^{4}}{M_{P}^{4}}+(2b_{p}+1-b_{k})\frac{m^{2}\phi^{2}}{2M_{P}^{4}}}{2\left[q^{4}(1+\zeta_{v})^{2}+(1-b_{p})\frac{\lambda}{4}\frac{\phi^{4}}{M_{P}^{4}}-(1+b_{p})\frac{m^{2}\phi^{2}}{2M_{P}^{4}}\right]},$ (2.36)
$K_{2}(\phi)=\frac{(1+b_{k})^{2}}{2\left[q^{4}(1+\zeta_{v})^{2}+(1-b_{p})\frac{\lambda}{4}\frac{\phi^{4}}{M_{P}^{4}}-(1+b_{p})\frac{m^{2}\phi^{2}}{2M_{P}^{4}}\right]}.$ (2.37)
Instead of the system of the Einstein equation (2.14) with the energy-momentum tensor (2.34)-(2.37) and the field $\phi$-equation obtained by substituting $\zeta$ from the constraint (2.17) into eq.(2.32), it is much more convenient to work with the TMT-effective action, the variation of which gives these equations. As usual, if $\tilde{g}^{\alpha\beta}\phi_{,\alpha}\phi_{,\beta}>0$, the energy-momentum tensor $T_{\mu\nu}^{(eff)}$ can be rewritten in the form of a perfect fluid. Then the pressure density plays the role of the matter Lagrangian in the effective action, and we arrive at the tree level TMT-effective action
$S_{eff}^{(TMT)}=\int\left(-\frac{M_{P}^{2}}{2}R(\tilde{g})+L_{eff}^{(tree)}(\phi,X_{\phi}))\right)\sqrt{-\tilde{g}}d^{4}x,$ (2.38)
with the following tree level TMT-effective Lagrangian
$L_{eff}^{(tree)}(\phi,X_{\phi})=K_{1}(\phi)\frac{X_{\phi}}{\Omega}-\frac{1}{2}K_{2}(\phi)\frac{X_{\phi}^{2}}{M_{P}^{4}}-U_{eff}^{(tree)}(\phi).$ (2.39)
Note that $L_{eff}^{(tree)}(\phi,X_{\phi}))$ has a form typical for K-essence models [40]-[43], and a manifestation of this is also the non-canonical form of the kinetic term in the $\phi$ equation (2.32). To extract the canonical kinetic term the equation must be divided by $\frac{b_{k}-\zeta}{(1+\zeta)\Omega}$. When studying the Higgs field equation in the vicinity of a vacuum in section 2.3, in the paragraph after eq.(2.22), we already encountered this feature, where, however, we had $\zeta=\zeta_{v}=const$ and $\Omega=1$.
In conclusion of this section, it can be stated that in the model under consideration, the equations and the TMT-effective (classical) action are applicable in the energy range from the electroweak energy scale (already considered in section 2.3) to the inflationary one, which we study in the next section.
3 The Two-Measure SM with small $\xi$ and Higgs inflation.
Preliminary consideration
3.1 About the shape of the TMT-effective potential and other features of the model
The model contains 6 parameters $V_{1}$, $V_{2}$, $b_{p}$, $b_{k}$, $\lambda$ and $m$. Instead of $V_{2}<0$ it is convenient to use parameter $q^{4}\leq 1$ defined by the relation (2.33). Instead of $V_{1}<0$ we use $\zeta_{v}$ defined by eq.(2.25). Further in this paper, for simplicity, we choose
$V_{1}=V_{2}\quad\Longleftrightarrow\quad\zeta_{v}=1.$ (3.1)
It will be clear from what follows that this choice is made without loss of generality. As a specific example of small $\xi$ from now on, we will always use $\xi=\frac{1}{6}$. We also choose
$b_{p}=\frac{1}{2}(1+\delta_{p})\quad 0<\delta_{p}\ll 1$ (3.2)
The impact of the choice $b_{p}\approx\frac{1}{2}$ and the value of $\delta_{p}$ will become clear in sections 4 and 5.2.
To apply the model to the study of the inflationary epoch, we would like to change field $\phi$ to the canonically normalized scalar field. For this to be done it is usually enough to redifine $\phi$ by means of the relation
$\phi=\frac{M_{P}}{\sqrt{\xi}}{\sinh}\left(\sqrt{\xi}\frac{\varphi}{M_{P}}\right).$ (3.3)
However, as can be seen from eq.(2.32), the redefinition (3.3) absorbs only the factor $\Omega^{-1}$, but does not provide canonical normalization of the field $\varphi$. The reason is that the TMT-effective action (2.38), (2.39) has the structure of K-essence type models. Thus, we are limited to only the redefinition (3.3) and arrive at the TMT-effective action in terms of $\varphi$
$S_{eff}^{(TMT)}=\int\left(-\frac{M_{P}^{2}}{2}R(\tilde{g})+L_{eff}^{(tree)}(\varphi,X_{\varphi})\right)\sqrt{-\tilde{g}}d^{4}x$ (3.4)
with the following tree-level TMT-effective Lagrangian rewritten in terms of $\varphi$ and $X_{\varphi}$
$L_{eff}^{(tree)}(\varphi,X_{\varphi})=K_{1}(\varphi)X_{\varphi}-\frac{1}{2}{\tilde{K}}_{2}(\varphi)\frac{X_{\varphi}^{2}}{M_{P}^{4}}-U_{eff}^{(tree)}(\varphi),$ (3.5)
where
$U_{eff}^{(tree)}(\varphi)=\frac{\lambda M_{P}^{4}}{4\xi^{2}}{\tanh}^{4}z\cdot F(z),$ (3.6)
$F(\varphi)=\frac{(\zeta_{v}+b_{p})q^{4}+\frac{\lambda}{16\xi^{2}}\sinh^{4}z-\frac{\lambda m^{2}}{4\xi M_{P}^{2}}\sinh^{2}z+\frac{m^{4}}{4\lambda M_{P}^{4}}-2(\zeta_{v}-b_{p})q^{4}\frac{\xi m^{2}}{\lambda M_{P}^{2}}\cdot\sinh^{-2}z}{(1+\zeta_{v})^{2}q^{4}+(1-b_{p})\frac{\lambda}{4\xi^{2}}{\sinh}^{4}z-(1+b_{p})\frac{m^{2}}{2M_{P}^{2}\xi}\sinh^{2}z}$ (3.7)
$z=\sqrt{\xi}\frac{\varphi}{M_{P}},\qquad X_{\varphi}=\frac{1}{2}\tilde{g}^{\alpha\beta}\varphi_{\alpha}\varphi_{\beta},$ (3.8)
$K_{1}(\varphi)=\frac{2q^{4}(1+\zeta_{v})(b_{k}-\zeta_{v})+(2b_{p}+b_{k}-1)\frac{\lambda}{4\xi^{2}}\sinh^{4}z+(2b_{p}+1-b_{k})\frac{m^{2}}{2M_{P}^{2}\xi}\sinh^{2}z}{2\Bigl{[}q^{4}(1+\zeta_{v})^{2}+(1-b_{p})\frac{\lambda}{4\xi^{2}}\sinh^{4}z-(1+b_{p})\frac{m^{2}}{2M_{P}^{2}\xi}\sinh^{2}z\Bigr{]}},$ (3.9)
${\tilde{K}}_{2}(\varphi)=\frac{(1+b_{k})^{2}\cosh^{4}z}{2\Bigl{[}q^{4}(1+\zeta_{v})^{2}+(1-b_{p})\frac{\lambda}{4\xi^{2}}\sinh^{4}z-(1+b_{p})\frac{m^{2}}{2M_{P}^{2}\xi}\sinh^{2}z\Bigr{]}}.$ (3.10)
The last expression for ${\tilde{K}}_{2}(\varphi)$ takes into account that $X_{\phi}=\cosh^{2}z\cdot X_{\varphi}$. The TMT-effective energy-momentum tensor given by eq.(2.34) now takes the following form
$T_{\mu\nu}^{(eff)}=K_{1}(\varphi)\cdot\big{(}\varphi_{,\mu}\varphi_{,\nu}-\tilde{g}_{\mu\nu}X_{\varphi}\big{)}-{\tilde{K}}_{2}(\varphi)\cdot\frac{X_{\varphi}}{M_{P}^{4}}\Big{(}\varphi_{,\mu}\varphi_{,\nu}-\frac{1}{2}\tilde{g}_{\mu\nu}X_{\phi}\Big{)}+\tilde{g}_{\mu\nu}U_{eff}^{(tree)}(\varphi).$ (3.11)
The constraint (2.17) represented through $\varphi$ becomes
$\zeta(\varphi,X_{\varphi})$
$=$ $\frac{2q^{4}\zeta_{v}(1+\zeta_{v})-(2b_{p}-1)\frac{\lambda}{4\xi^{2}}\sinh^{4}z-(2b_{p}+1)\frac{m^{2}}{2\xi M_{P}^{2}}\sinh^{2}z+(1+b_{k})\cosh^{4}z\cdot\frac{X_{\varphi}}{M_{P}^{4}}}{2q^{4}(1+\zeta_{v})+\frac{\lambda}{4\xi^{2}}\sinh^{4}z-\frac{m^{2}}{2\xi M_{P}^{2}}\sinh^{2}z-(1+b_{k})\cosh^{4}z\cdot\frac{X_{\varphi}}{M_{P}^{4}}},$
Subsequently, it will be shown that in the process of inflation, which occurs in the slow-roll regime, the difference from the case of a canonically normalized scalar field turns out to be insignificant.
Despite the fact that in the model under study the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$ has a plateau-like character, there are two significant circumstances, without a detailed analysis of which it is impossible to obtain sufficiently definite quantitative predictions for comparing them with observational cosmological data. The first of them is the difference in the form of the function $U_{eff}^{(tree)}(\varphi)$ from those known in the literature. The second is due to the fact that TMT-effective action has a K-essence type structure. Without an adequate response to these problems, it would be impossible, for example, to ensure that inflation proceeds in a slow-roll regime. As will be shown in the following sections, there is a fairly wide range of model parameters in which the influence of the K-essence structure on the predicted inflation parameters turns out to be extremely insignificant. A noticeable difference from the predictions of a model with the same potential, but with a standard, non-K-essence structure can only exist at the end of inflation, which will be discussed in section 6.3. Below, in this section, we will completely ignore the K-essence structure, and therefore the results obtained should be considered as preliminary.
To get an idea of the shape of the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$, we note that its formula differs by the factor $F(\varphi)$, eq.(3.7), from the corresponding formulas for the potential in the T-model with n = 2 in the notations of ref. [28] and for the form of the Eistein-frame potential of the Palatini inflation models [12] (see also reviews [18], [19] and references therein). It is well known that to ensure agreement with Planck’s data, the height of the potential plateau should be of the order $\sim 10^{-10}M_{P}^{4}$. Therefore, depending on the relationship between parameters $q^{4}$ and $\frac{\lambda}{4\xi^{2}}$ two types of TMT potential shape are possible: with one plateau and with two plateaus. An intermediate form is also possible, when the presence of a second plateau is weakly expressed, but we will consider only cases with clearly defined one or two plateaus. When $\varphi>\frac{M_{P}}{\sqrt{\xi}}$ is such that numerator and denominator in $F(\varphi)$ are dominated by the terms $\propto\sinh^{4}z$, then $F(\varphi)\approx\frac{1}{4(1-b_{p})}$ and the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$ has one plateau
$U_{eff}^{(tree)}(\varphi)\Big{|}_{\varphi\gg\frac{M_{P}}{\sqrt{\xi}}}=\frac{\lambda M_{P}^{4}}{16(1-b_{p})\xi^{2}}{\tanh}^{4}\Bigl{(}\sqrt{\xi}\frac{\varphi}{M_{P}}\Bigr{)}$ (3.13)
However, generically it is possible that for lower values of $\varphi>\frac{M_{P}}{\sqrt{\xi}}$, such that $(\zeta_{v}+b_{p})q^{4}>\frac{\lambda}{16\xi^{2}}\sinh^{4}\Bigl{(}\sqrt{\xi}\frac{\varphi}{M_{P}}\Bigr{)}$ and $(1+\zeta_{v})^{2}q^{4}>(1-b_{p})$, there is a second plateau [^8] A similar situation can also be seen in ref. [35] described mainly by
$U_{eff}^{(tree)}(\varphi)\Big{|}_{\varphi>\frac{M_{P}}{\sqrt{\xi}}}=\frac{\lambda(\zeta_{v}+b_{p})M_{P}^{4}}{4(1+\zeta_{v})^{2}\xi^{2}}{\tanh}^{4}\Bigl{(}\sqrt{\xi}\frac{\varphi}{M_{P}}\Bigr{)}.$ (3.14)
A typical shape of the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$ with two plateaus is shown in figure 1. Apparently the second plateau must be responsible for the last stage of inflation, and the potential with two plateaus may be of interest as another type of inflation pattern. However, in this paper we will not consider this possibility in detail, but will focus on studying the model in the parameter range, where the effect of the existence of a second potential plateau does not manifest itself explicitly.
Figure 1: The graph of the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$, defined by eqs.(3.6) and (3.7) for the model parameters $\lambda=2.3\cdot 10^{-11}$,   $\xi=\frac{1}{6}$,   $m=0.7GeV$,   $b_{p}\approx 0.5$ and $q^{4}=3\cdot 10^{-4}$. This choice of $q^{4}$ means that $|V_{1}|=|V_{2}|\approx(10^{16}GeV)^{4}$, and this leads to a potential with two clearly defined plateaus (see also eqs.(3.13) and (3.14))
Figure 2: The curve labeled 1 is a plot of the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$, defined by eqs.(3.6) and (3.7), has one plateau due to the choice of $q^{4}=3\cdot 10^{-10}$, i.e. $|V_{1}|=|V_{2}|\approx(10^{16}GeV)^{4}$, while other model parameters are the same as in Fig. 1. The curves labeled 2 and 3 are plots of the scalar function $\zeta(\varphi)$ defined by the constraint (3.1), where the terms $\propto\frac{X_{\varphi}}{M_{P}^{4}}$ are assumed to be negligible compared to the other terms. Curve 2 is the plot of $\zeta(\varphi)$ if the model parameter $b_{p}=0.55$ is chosen, and the plot intersects $\zeta=0$ at $\varphi=4.25M_{P}$. Curve 3 is the plot of $\zeta(\varphi)$ if the model parameter $b_{p}=0.5(1+10^{-8})$ is chosen, and the plot intersects $\zeta=0$ at $\varphi\approx 14M_{P}$.
  • The TMT-effective potential with one plateau
    Let us now analyze in more detail the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$ with one plateau. With the choice described by eqs.(3.1) and (3.2), and using the parameter values $\lambda=2.3\cdot 10^{-11}$,   $\xi=\frac{1}{6}$,   $m=0.7GeV$, one can see that the TMT-effective potential with one plateau can be achieved [^9] For $\varphi>\frac{M_{P}}{\sqrt{\xi}}=\sqrt{6}M_{P}$, we find that $F(\varphi)\approx\frac{1}{2}$ with a relative error $\frac{|F(\varphi)-\frac{1}{2}|}{1/2}<2.2\cdot 10^{-3}$. if $q^{4}\lesssim 10^{-7}$. For further analysis of the model, we choose
    $q^{4}=3\cdot 10^{-10},\quad\text{that is}\quad V_{1}=V_{2}\approx-(10^{16}GeV)^{4}.$ (3.15)
    The graph of the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$, given by eqs.(3.6) and (3.7), is shown in figure 2. Substituting $\xi=\frac{1}{6}$ and $b_{p}=\frac{1}{2}$ into (3.13), we find that for $\varphi>6M_{P}$ the TMT-effective potential is described with high accuracy by the formula
    $U_{eff}^{(tree)}(\varphi)|_{\varphi>6M_{P}}=\frac{\lambda M_{P}^{4}}{8\xi^{2}}{\tanh}^{4}\Bigl{(}\frac{\varphi}{\sqrt{6}M_{P}}\Bigr{)}\approx\frac{\lambda M_{P}^{4}}{8\xi^{2}}\Bigl{(}1-8e^{-\sqrt{\frac{2}{3}}\frac{\varphi}{M_{P}}}\Bigr{)}.$ (3.16)
    Note that due to $F(\varphi)\approx\frac{1}{2}$ at $\varphi>6M_{P}$, the formula for the TMT-effective potential (3.16) differs from the form of the potential in the Einstein frame in the Palatini-Higgs models [12], [18], [19] by a factor $\frac{1}{2}$. But the fundamental difference from the Palatini-Higgs models is that here $\xi=\frac{1}{6}$. In this respect, the potential of the T-model [28] is more suitable for comparison: the TMT-effective potential $U_{eff}^{(tree)}(\varphi)|_{\varphi>6M_{P}}$ coincides with the T-model potential in the region $\varphi>6M_{P}$ for the case n=2 in the notation of the paper [28].
3.2 Higgs inflation. Preliminary consideration of the slow-roll approximation
Continuing our preliminary consideration, let us start with the $\varphi$ field equation and the Friedmann equation in the spatially flat FLRW Universe completely ignoring the K-essence type structure of the TMT-effective action (3.4), (3.5)
$\ddot{\varphi}+3H\dot{\varphi}+U_{eff}^{(tree)\,\prime}=0,$ (3.17)
$H^{2}=\frac{1}{3M_{P}^{2}}\Bigl{(}\frac{1}{2}\dot{\varphi}^{2}+U_{eff}^{(tree)}(\varphi)\Bigr{)}$ (3.18)
An overdot denotes $d/dt$ and a prime denotes $d/d\varphi$. In slow-roll approximation $\frac{1}{2}\dot{\varphi}^{2}\ll U_{eff}$ and $|\ddot{\varphi}|\ll|U_{eff}^{(tree)\,\prime}|\approx 3H|\dot{\varphi}|$ these equations reduce to
$3H\dot{\varphi}\simeq U_{eff}^{(tree)\,\prime}(\varphi),$ (3.19)
$H^{2}\simeq\frac{1}{3M_{P}^{2}}U_{eff}^{(tree)}(\varphi),$ (3.20)
from where it follows that the flatness conditions imposed on $U_{eff}^{(tree)}(\varphi)$
$\epsilon=\frac{M_{P}^{2}}{2}\Biggl{(}\frac{U_{eff}^{(tree)\,\prime}}{U_{eff}^{(tree)}}\Biggr{)}^{2}\ll 1,\qquad|\eta|=M_{P}^{2}\Biggl{|}\frac{U_{eff}^{(tree)\,\prime\prime}}{U_{eff}}\Biggr{|}\ll 1$ (3.21)
must be satisfied. In sec.6 we will study in detail the modification of the flatness conditions if the K-essence type structure is taken into account.
At this stage of the study, our main goal is to understand how Higgs inflation with a small model parameter $\xi$ of nonminimal coupling becomes possible in the TMSM. First of all, choosing the value of $\xi=\frac{1}{6}$ we must find the value of the primordial parameter $\lambda$ in such a way that the model predictions agree with the CMB data. In light of what was said at the very end of the previous subsection, the predictions for the values of all perturbation parameters in the model under study must coincide with the corresponding predictions of the T-model. Therefore, to find $\lambda$, it is sufficient, for example, to consider the prediction for scalar perturbations $\mathcal{A}_{s}$. Using the standard method for estimating the duration of inflation, measured by the number of e-folds $N$, we obtain
$\epsilon\approx\frac{1}{8\xi N^{2}}$ (3.22)
The power spectrum of scalar perturbations $\mathcal{A}_{s}$, measured in the CMB at large scales, can be expressed [54] in terms of the slow-roll parameter $\epsilon$ as $\mathcal{A}_{s}\approx\frac{1}{24\pi^{2}\epsilon}\frac{U_{eff}}{M_{P}^{4}}$, and for large $N$ we get
$\mathcal{A}_{s}\approx\frac{\lambda N^{2}}{24\pi^{2}\xi}.$ (3.23)
With the observational data bound [20], [21] $\mathcal{A}_{s}\approx 2.1\cdot 10^{-9}$ we get the relation
$\lambda\approx 5\cdot 10^{-7}\frac{\xi}{N^{2}}.$ (3.24)
For $N=60$, with our choice of parameter $\xi=\frac{1}{6}$, we obtain
$\lambda\approx 2.3\cdot 10^{-11}.$ (3.25)
Now, staying within the framework of a preliminary study, we will take the final step to demonstrate how the TMSM model proposed in this paper allows us to implement the Higgs inflation model with $\xi=\frac{1}{6}$ while satisfying the observed cosmological data due to $\lambda\sim 10^{-11}$, on the one hand, and to satisfy the requirement of $\lambda\sim 0.1$ near vacuum imposed by the SM, on the other hand. This is possible due to the arbitrariness of the primordial parameter $\lambda$ in TMSM, as discussed at the end of section 2.3. Indeed, as can be seen from eqs.(2.30) and (2.31), to obtain the desired VEV it suffices to choose the model parameter $b_{k}$ such that the smallness of $b_{k}-\zeta_{v}=b_{k}-1$ compensates for the smallness of $\lambda$. But as already noted in footnote 5, in ref. [22] it will be shown that the VEV of the Higgs field, which should be compared with the VEV in the SM, differs from the value of $\sigma$ given by eq.(2.31). However, in this paper, the exact value of $b_{k}-1$ does not play a significant role. Therefore, here it is sufficient to restrict ourselves to the order of magnitude of the result obtained in ref. [22]: $b_{k}-\zeta_{v}\sim 10^{-5}$. Thus, our goal is achieved by simply choosing a value for the parameter $b_{k}$ very close to $\zeta_{v}=1$. Turning to how the primordial action (2.1), (2.2) is structured, we understand that such a choice seems quite natural. Indeed, the extremely small deviation of $b_{k}$ from unity can be interpreted as a quantum effect of a more fundamental theory.
4 More about the $\zeta(\phi)$ function and the TMT-effective potential.
TMSM’s answer to the mystery of the Higgs potential structure
With the obtained value of the parameter $\lambda$, eq.(3.25), we return to the discussion of the shape of the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$, eqs.(3.6) and (3.7), the graph of which in the entire region $\varphi\geq 0$ is shown [^10] Remember that we chose $q^{4}=3\cdot 10^{-10}$ which ensures the TMT potential with one plateau. in figure 2. Two graphs of $\zeta(\varphi)$ for the case of $X_{\varphi}=0$ are also presented in figure 2; for these two graphs, in addition to the parameters listed above, two corresponding values of $\delta_{p}=0.1$ and $\delta_{p}=10^{-8}$ are chosen in the definition $b_{p}=0.5(1+\delta_{p})$ given by eq.(3.2). As will be explained in section 5, for any admissible values of $\varphi$ the function $\zeta$ must be positive. Therefore, from the shape of curve 2 in figure 2 it follows that the choice $\delta_{p}=0.1$ is unsatisfactory. It should be noted that the graph of $\zeta(\varphi)$ for $\delta_{p}=10^{-8}$, represented by curve 3 in figure 2, is to some extent illustrative, since the case $X_{\varphi}=0$ is unrealistic in general. However, as shown in ref. [22], in the regime of slow-roll inflation, which occurs on the plateau part of $U_{eff}^{(tree)}(\varphi)$, i.e., for $\varphi>6M_{P}$, the contribution of $X_{\varphi}$ to the constraint (3.1) is bounded from above by the following estimate $\frac{X_{\varphi}}{M_{P}^{4}}\approx 128\lambda\cdot e^{-2\sqrt{\frac{8}{3}}\frac{\varphi}{M_{P}}}<9\cdot 10^{-18}$. Therefore, for $\varphi>6M_{P}$, the curve 3 in figure 2 can be considered as quite reliable graph for $\zeta$. In the opposite region, where $\varphi\lll M_{P}$, one can expect that $X_{\varphi}\lll M_{P}^{4}$ and therefore the $\zeta(\varphi)$ graph is also reliable. However, it is quite natural to suppose that soon after the end of inflation the contribution of $X_{\varphi}$ to the constraint (3.1) may become significant. Nevertheless, from the graph with $\delta_{p}=10^{-8}$ it follows that the decrease in $\varphi$ at the inflationary stage is accompanied by a monotonous increase in $\zeta(\varphi)$ from positive values close to zero [^11] The fundamental significance of the condition $\zeta>0$ will be the subject of study in section 5.; and for $\varphi\lll M_{P}$, with a decrease in $\varphi$, $\zeta(\varphi)$ also increases monotonically, asymptotically approaching the value $\zeta=\zeta_{v}=1$.
The representation of $U_{eff}^{(tree)}(\varphi)$ in figure 2 does not allow to see important details of the shape, which differ greatly in the regions $\varphi>M_{P}$ and $\varphi\ll M_{P}$. This is due to the greatly different behaviour of the factor $F(\varphi)$, eq.(3.7), in the regions $\varphi>M_{P}$ and $\varphi\ll M_{P}$. The analysis simplified if we use the value of the primordial model parameter $m\approx 0.7GeV$ obtained in ref. [22], that is, $\frac{m^{2}}{M_{P}^{2}}\approx 8\cdot 10^{-38}$. Then the form of the function $F(\varphi)$ depends on the ratio of the second and fifth terms in the numerator of $F(\varphi)$. A direct estimate shows that the value $\varphi\approx 1.7\cdot 10^{-4}M_{P}\approx 4.1\cdot 10^{14}GeV$ divides the entire region $\varphi\geq 0$ into two intervals, where one of the indicated terms dominates over the other. Accordingly, we arrive at the following asymptotic representation for the TMT-effective potential in two opposite regions
$U_{eff}^{(tree)}(\varphi)=\begin{cases}\frac{\lambda M_{P}^{4}}{8\xi^{2}}{\tanh}^{4}z\cdot\frac{3q^{4}+\frac{\lambda}{8\xi^{2}}\sinh^{4}z}{4q^{4}+\frac{\lambda}{8\xi^{2}}\sinh^{4}z}&\text{for}\quad\varphi\gg 10^{-4}M_{P}\\ \frac{\zeta_{v}+b_{p}}{(1+\zeta_{v})^{2}}\frac{\lambda}{4}\phi^{4}-\frac{\zeta_{v}-b_{p}}{(1+\zeta_{v})^{2}}\frac{m^{2}}{2}\phi^{2}&\text{for}\quad\phi\ll 10^{-4}M_{P},\end{cases}$ (4.1)
where we also took into account that $\varphi\approx\phi$ when $\varphi\ll M_{P}$.
The result in region $\phi\ll 10^{-4}M_{P}$, presented in figure 3, will be more understandable if we consider the expression for the TMT-effective Lagrangian $L_{eff}^{(tree)}(\varphi,X_{\varphi})$ given by eqs.(3.4)-(3.10). Considering that, as follows from eqs.(3.9) and (3.10)
$K_{1}|_{\phi\ll 10^{-4}M_{P}}=\frac{b_{k}-\zeta_{v}}{1+\zeta_{v}},\quad{\tilde{K}}_{2}|_{\phi\ll 10^{-4}M_{P}}=\frac{(1+b_{k})^{2}}{2q^{4}(1+\zeta_{v})^{2}},$ (4.2)
we arrive at the following expression
$L_{eff}^{(tree)}|_{\phi\ll 10^{-4}M_{P}}=\frac{b_{k}-\zeta_{v}}{1+\zeta_{v}}X_{\phi}-\frac{(1+b_{k})^{2}}{4q^{4}(1+\zeta_{v})^{2}}\frac{X_{\phi}}{M_{P}^{4}}X_{\phi}-U_{eff}^{(tree)}(\phi)|_{\phi\ll 10^{-4}M_{P}}.$ (4.3)
Figure 3: The plot of the TMT-effective potential $U_{eff}^{(tree)}(\phi)$ for $\phi\ll 10^{-4}M_{P}$, defined in eq.(4.1), has the well-known form of the Higgs potential. In the $\phi$-equation of motion with the canonical kinetic term, the derivative of this potential turns out to be divided by a constant factor $\frac{b_{k}-\zeta_{v}}{1+\zeta_{v}}$, where $b_{k}-\zeta_{v}=b_{k}-1\approx 10^{-5}$.
In the region $\phi\ll 10^{-4}M_{P}$ it is natural to assume that $|X_{\phi}|\lesssim|U_{eff}(\phi)|_{\phi\ll 10^{-4}M_{P}}|\ll q^{4}M_{P}^{4}\sim(10^{16}GeV)^{4}$. Then it is easy to see that the second term in (4.3) is many orders of magnitude smaller than the first. In this approximation, the field $\phi$ equation of motion, obtained using the Lagrangian $L_{eff}^{(tree)}|_{\phi\ll 10^{-4}M_{P}}$ and then dividing by a constant factor $\frac{b_{k}-\zeta_{v}}{1+\zeta_{v}}$, coincides with the $\phi$-equation near the vacuum, described by eqs.(2.29) and (2.30), found in section 2.3. This analysis allows us to get some idea of what structure the field equations might have for $\phi\lesssim 10^{-4}M_{P}$.
  • We can now understand how in the TMSM, despite the fact that the parameter $m^{2}$ in the primordial action is positive, the mass term near the vacuum turns out to have the "wrong" sign, which provides a spontaneous symmetry breaking. This mystery of the Higgs potential is explained in the TMSM as a consequence of the behavior of the function $\zeta(\varphi)$ during cosmological evolution, which was discussed at the beginning of this section. This can be best seen by analyzing the $\phi$-equation in the $\zeta$-dependent form (2.19) (which also has a non-kanonical kinetic term). In the process of the increase of the function $\zeta(\varphi)$ described above, at a certain moment of cosmological evolution $\zeta(\varphi)$ crosses the value $b_{p}\approx\frac{1}{2}$; as a result, the mass term proportional to $(b_{p}-\zeta)m^{2}\phi$ in eq.(2.19) changes sign to the "wrong" one.
5 Constraint, orientability of the space-time manifold
and their impact on cosmology
The question of why the initial conditions for the solution of cosmological equations ensure the occurrence of inflation is of fundamental theoretical importance and is known in the literature [23], [24], [30]-[34] as the problem of the initial conditions for inflation. In this section we will show that in the model under study the conditions described by inequalities (1.1) are guaranteed to be met. However, in section 6 we will discover that the effects of the TMT-effective K-essence type structure lead to the need to take into account that, instead of inflation, cosmological evolution may begin with some exotic scenario obeying completely different laws of physics. For this reason, we are forced in this section to formulate the problem and the method existing in TMSM for solving it in a broader context. Namely, if the solutions of cosmological equations describe the evolution of the universe (in classical space-time), then: 1) the basic structure of the TMT, which also underlies the TMSM, imposes dynamics-independent constraints on the admissible initial conditions of solutions; 2) in the following section 6 we will show that the dynamics features in the model under study impose stronger constraints on the admissible initial conditions than those mentioned in item 1).
5.1 Brief summary of ref. [35]: guaranteed initial conditions for inflation
and the Borde-Guth-Vilenkin theorem
A solution (2.8) of eq.(2.4), the general form of which is the same as in ref. [35], exists under the condition $\Upsilon(x)\neq 0$, Eq.(2.7), i.e, only if $\Upsilon(x)>0$ or only if $\Upsilon(x)<0$. Therefore, only those solutions of the system of equations are valid for which the corresponding $\Upsilon(x)$ is sign-definite, since Eq.(2.4) is one of the equations resulting from the principle of least action. It should be noted here that, as usual, by default we assume that the original metric $g_{\mu\nu}$ in the primordial action is regular, that is $g=\det(g_{\mu\nu})<0$. Therefore, the validity of the solutions regarding the fulfillment of the condition on the sign of $\Upsilon$ can be controlled by checking the sign of the scalar $\zeta=\frac{\Upsilon}{\sqrt{-g}}$. In particular, in the model under consideration, with our choice of parameters and constant of integration, the value of $\zeta$ in vacuum, eq.(2.25), is positive: $\zeta_{v}>0$. Thus, only those cosmological solutions (together with their initial conditions) are valid for which $\zeta(x)$ is positive throughout the evolution of the universe. In other words, these solutions lose their validity when we try to extend them to the region where $\zeta$ crosses zero and becomes negative. The condition $\Upsilon(x)>0$ is equivalent to the sign-definitness of the volume 4-form $\Upsilon(x)d^{4}x$ on the 4D-space-time manifold. From the geometrical point of view this means that the differential manifold of our Universe is orientable [^12] For more mathematical aspects of this issue see, e.g., books [55], [56] and section 4 in ref. [35]..
If $\zeta(x)$ vanishes for a certain field configuration, then this situation is of particular interest and was studied in detail in paper [35] within a fairly simple model. As a result of the condition $\Upsilon(x)>0$, or equivalently $\zeta(x)>0$, we discovered that, if $X_{\varphi}>0$, there exists an upper bound $\varphi_{0}$ on admissible values of the inflaton field $\varphi$ for which
$\Upsilon\equiv\zeta(\varphi,X_{\varphi})\cdot\sqrt{-g}\rightarrow 0^{+}\quad{\text{a}nd}\quad X_{\varphi}\rightarrow 0^{+}\quad{\text{a}s}\quad\varphi\rightarrow\varphi_{0}^{\,\,\,-}.$ (5.1)
A detailed study has established that if $X_{\varphi}^{(in)}>0$, that is the initial kinetic energy density $\rho_{kin}^{(in)}$ prevails over initial gradient energy density $\rho_{grad}^{(in)}$, then there is an interval of initial values $\varphi_{in}^{(min)}\leq\varphi_{in}<\varphi_{0}$, where $\rho_{kin}^{(in)}$ and $\rho_{grad}^{(in)}$ cannot exceed the potential energy density and hence the initial conditions necessary for the onset of infation are satisfied. In addition, it was shown that $\Upsilon(x)=0$ is a spacelike hypersurface such that in the inflating spacetime of the cosmological model studied in ref. [35], timelike geodesics cannot be extended to the past beyond $\Upsilon(x)=0$. This conclusion is consistent with the statement of the BGV theorem [38], and we had reason to assert that the hypersurface $\Upsilon(x)=0$ is the boundary $\mathcal{B}$, which appears in the formulation of the theorem. The analyze made in ref. [35] shows that $\mathcal{B}$ is the boundary of the space-time manifold of our Universe. Here we will limit ourselves to this brief summary of the essence of the discovered effect and move on to the analysis of the constraint, its features and consequences for the Higgs inflation model studied in this paper.
5.2 Constraint, initial conditions
and orientability of the space-time manifold
The constraint (3.1) has a fundamental role throughout the entire process of cosmological evolution, but we will consider here only the case of $\sqrt{\xi}\varphi>M_{P}$ applicable to the inflationary stage. We are going to show that the condition $\zeta>0$ imposes restrictions on the possible values of $\varphi$ and $X_{\varphi}$ of the solutions to the cosmological equations. In the context of this paper, the most important question for us is to find out whether the application of these restrictions to the initial values $\varphi_{in}$ and $X_{\varphi}^{(in)}$ ensures the onset of inflation. However, the need to take into account the influence of the K-essence type structure forces us to limit ourselves here to finding the quantitative restrictions imposed on the initial values $\varphi_{in}$ and $X_{\varphi}^{(in)}$ by the condition $\zeta>0$, and to postpone the study of their role in the onset of inflation until section 6.1.
It turns out that the signs with which the terms with $X_{\varphi}$ enter into the constraint (3.1) are opposite to the signs of the corresponding terms in the model of ref. [35]. This difference leads to restrictions on the possible values of $\varphi$ and $X_{\varphi}$ of the solutions to the cosmological equations, which differ significantly from the model results of ref. [35]. Therefore, the restrictions on the initial values $\varphi_{in}$ and $X_{\varphi}^{(in)}$ of these solutions also differ significantly from the model results of ref. [35]. For $\sqrt{\xi}\varphi>M_{P}$ the constraint (3.1)) can be conveniently rewritten by omitting terms $\propto\frac{m^{2}}{M_{P}^{2}}$
$\zeta(\varphi,X_{\varphi})=\frac{32q^{4}e^{-\sqrt{\frac{8}{3}}\frac{\varphi}{M_{P}}}-\frac{\lambda}{8\xi^{2}}\delta_{p}+\frac{X_{\varphi}}{M_{P}^{4}}}{32q^{4}e^{-\sqrt{\frac{8}{3}}\frac{\varphi}{M_{P}}}+\frac{\lambda}{8\xi^{2}}-\frac{X_{\varphi}}{M_{P}^{4}}},$ (5.2)
where we took $\xi=\frac{1}{6}$, $\zeta_{v}=1$, $b_{k}=1$ and the presentation (3.2) of $b_{p}$ via $\delta_{p}$ was used. It is obvious that $\zeta(\varphi,X_{\varphi})>0$ if and only if the values of $\varphi$ and $X_{\varphi}$ satisfy the following double inequality
$\frac{\lambda}{8\xi^{2}}\delta_{p}-32q^{4}e^{-\sqrt{\frac{8}{3}}\frac{\varphi}{M_{P}}}<\frac{X_{\varphi}}{M_{P}^{4}}<\frac{\lambda}{8\xi^{2}}+32q^{4}e^{-\sqrt{\frac{8}{3}}\frac{\varphi}{M_{P}}}.$ (5.3)
It is important to note that the way this double inequality is derived has nothing to do with the K-essence type structure. Both the solutions to the cosmological equations obtained from the TMT-effective action (3.4)-(3.10) and the initial conditions of these solutions must satisfy the requirements imposed by this double inequality. Therefore, they must be fulfilled throughout the entire process of cosmological evolution, driven by the classical scalar field $\varphi$ (under condition $\sqrt{\xi}\varphi>M_{P}$), including the very beginning of evolution. Consequently, by applying the inequalities for the initial values [^13] It should be kept in mind that $\varphi_{in}$ and $X_{\varphi}^{(in)}$ are treated as independent of each other. $\varphi_{in}$ and $X_{\varphi}^{(in)}$, we will be able to see what restrictions these inequalities impose on the initial conditions, which is of particular interest in the light of clarifying the issue of the occurrence of inflation.
To interpret the results, one must keep in mind that $X_{\varphi}$ equals to the difference of the canonical [^14] It should be noted that only the regular form of kinetic and gradient energy densities is used here, and the kinetic and gradient energy densities modified by the K-essence type structure will be considered in section 6. kinetic and gradient energy densities
$X_{\varphi}=\frac{1}{2}\tilde{g}^{\alpha\beta}\varphi_{,\alpha}\varphi_{,\beta}=\frac{1}{2}\left(\dot{\varphi}^{2}-\frac{1}{a^{2}}(\nabla\varphi)^{2}\right)=\rho_{kin}-\rho_{grad},$ (5.4)
Therefore, the initial $X_{\varphi}^{(in)}=\rho_{kin}^{(in)}-\rho_{grad}^{(in)}$ can be positive or negative depending on how inhomogeneous and anisotropic at the beginning of classical evolution was the space domain whose expansion generates our Universe.
Let’s look at the implications that can be drawn from the inequalities (5.3) with the model parameters used in figure 2, where we chosed $q^{4}=3\cdot 10^{-10}$. The shape of the potential $U_{eff}^{(tree)}(\varphi)$ shows that, with the exception of the very end, inflation can occur at $\varphi\gtrapprox 6M_{P}$; if evolution starts from other scenario, possibility of which will be found in sec.6, it is natural to assume that this also occurs at $\varphi\gtrapprox 6M_{P}$. The height of the potential plateau is between $10^{-10}M_{P}^{4}\lesssim U_{eff}(\varphi)\lesssim 1.04\cdot 10^{-10}M_{P}^{4}$. The latter will be used in the numerical estimates made below. The remaining free parameter is $10<\delta_{p}<$ 1 and it turns out that changing its value can lead to qualitatively different results. As numerical examples, consider two cases $\delta_{p}=0.1$ and $\delta_{p}=10^{-8}$, corresponding to two graphs $\zeta(\varphi)$ for $X_{\varphi}=0$ in figure 2.
  • In the case $\delta_{p}=0.1$, that is $b_{p}=0.55$, the inequalities read
    $10^{-11}\lesssim\frac{X_{\varphi}}{M_{P}^{4}}\lesssim\frac{\lambda}{8\xi^{2}}\approx\frac{U_{eff}^{(tree)}(\varphi)}{M_{P}^{4}}\approx 10^{-10}.$ (5.5)
    In the case of a homogeneous and isotropic universe, where $X_{\varphi}=\frac{1}{2}\dot{\varphi}^{2}$, we see that $\dot{\varphi}^{2}$ is limited to positive values. There are at least two reasons why this result is unacceptable. First, the existence of this lower positive limit for $X_{\varphi}$ would lead to the completely unnatural conclusion that at the very beginning of cosmological evolution in classical space-time the existence of regions in which $\frac{1}{a^{2}}(\nabla\varphi)^{2}>\frac{1}{2}\dot{\varphi}^{2}$ is impossible. Second, note that, as follows from the plots of $\zeta(\varphi)$ for $X_{\varphi}=0$ in Fig.2, in the case of $\delta_{p}=0.1$, values of $X_{\varphi}>0$ close to zero are prohibited for $\varphi>4.25.M_{P}$, since in this case $\zeta<0$. All of the above makes the choice of the parameter $b_{p}=0.55$ undesirable.
  • In the case
    $b_{p}=0.5(1+10^{-8}),$ (5.6)
    for $\varphi>6M_{P}$, the inequalities (5.3) are reduced to
    $-5.3\cdot 10^{-13}\lesssim\frac{X_{\varphi}}{M_{P}^{4}}\lesssim\frac{\lambda}{8\xi^{2}}\approx\frac{U_{eff}^{(tree)}(\varphi)}{M_{P}^{4}}\approx 10^{-10}$ (5.7)
    with very high accuracy. Obviously, in this example, at the very beginning of cosmological evolution in classical space-time, the existence of regions in which $\frac{1}{a^{2}}(\nabla\varphi)^{2}>\frac{1}{2}\dot{\varphi}^{2}$ is not forbidden. For initial values of $\varphi_{in}$ and $(\partial_{\alpha}\varphi)_{in}$ we obtain
    $-5.3\cdot 10^{-13}M_{P}^{4}\lesssim\rho_{kin}^{(in)}-\rho_{grad}^{(in)}\lesssim U_{eff}^{(tree)}(\varphi_{(in)})\approx 10^{-10}M_{P}^{4}$ (5.8)
    Discussing the possible sign of $X_{\varphi}^{(in)}=\rho_{kin}^{(in)}-\rho_{grad}^{(in)}$, it cannot be ruled out that $\rho_{kin}^{(in)}>U_{eff}^{(tree)}(\varphi_{in})$ and $\rho_{grad}^{(in)}>U_{eff}^{(tree)}(\varphi_{in})$ , but $X_{\varphi}^{(in)}<U_{eff}^{(tree)}(\varphi_{in})$. However, this possibility seems unlikely, and we will not consider it. Taking into account this remark, in the case where $\rho_{kin}^{(in)}>\rho_{grad}^{(in)}$, the inequality (5.8) reduces to
    $\frac{1}{2}\dot{\varphi}_{in}^{2}\lesssim U_{eff}^{(tree)}(\varphi_{(in)})\approx 10^{-10}M_{P}^{4};$ (5.9)
    and if $\rho_{grad}^{(in)}>\rho_{kin}^{(in)}$ then the inequality (5.8) reduces to
    $\frac{1}{a^{2}}(\nabla\varphi)_{in}^{2}\lesssim 5.3\cdot\cdot 10^{-13}M_{P}^{4}\ll U_{eff}^{(tree)}(\varphi_{in}).$ (5.10)
Thus, we come to the conclusion that there is a quite wide range of model parameters in which conditions (1.1) are satisfied. At this stage, it is worth noting once again that the described results follow only from the condition $\zeta>0$, which, as discussed above, means that the space-time manifold of our Universe is orientable. Recall also that, as explained in the paragraph after eq.(2.17), the origin of the restrictions imposed by the constraint on the dynamics of the field $\phi$ are in no way dependent on the dynamics of the field $\phi$; this is true in general and also in this particular case with respect to the condition $\zeta>0$. However, as noted at the beginning of this section, in order to assert that the conditions necessary for the onset of inflation are met, we must take into account the influence of the K-essence type structure.
6 Impact of the K-essence type structure
on the first stages of classical cosmological evolution
6.1 General overview and possible scenarios
for the beginning of cosmological evolution
Using the TMT-effective action, eqs.(3.4)-(3.10), in the spatially flat FLRW universe, the following Friedmann and scalar field equations are obtained after using eqs.(3.11), (3.4), (3.5):
$H^{2}=\frac{1}{3M_{P}^{2}}\Bigl{[}\Big{(}K_{1}-\frac{3}{2}\tilde{K}_{2}\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\Big{)}\cdot\frac{1}{2}\dot{\varphi}^{2}+U_{eff}^{(tree)}(\varphi)\Bigr{]},$ (6.1)
$\Bigl{(}K_{1}-3\tilde{K}_{2}\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\Bigr{)}\ddot{\varphi}+\Bigl{(}K_{1}-\tilde{K}_{2}\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\Bigr{)}3H\dot{\varphi}+\Bigl{(}K_{1}^{\prime}-\tilde{K}^{\prime}_{2}\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\Bigr{)}\dot{\varphi}^{2}+U_{eff}^{(tree)^{\prime}}(\varphi)=0.$ (6.2)
We will continue to explore the model by adding to the parameters used in figure 2, parameter $b_{k}=1+\mathcal{O}(10^{-5})$. The graphs of $K_{1}(\varphi)$ and $\tilde{K}_{2}(\varphi)$ are presented in figure 4. Except at the end, in this section we are only interested in the region of $\varphi$, where $\sqrt{\frac{8}{3}}\varphi\gg M_{P}$. In this region, the analytical form of $K_{1}(\varphi)$ and $\tilde{K}_{2}(\varphi)$ reduces to
$K_{1}(\varphi)\approx 1-185.5\cdot e^{-\sqrt{\frac{8}{3}}\frac{\varphi}{M_{P}}},$ (6.3)
$\tilde{K}_{2}(\varphi)\approx 1.93\cdot 10^{10}\Big{(}1-185.5\cdot e^{-\sqrt{\frac{8}{3}}\frac{\varphi}{M_{P}}}\Bigr{)}.$ (6.4)
In the region $\varphi>6M_{P}$, where the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$ is flat with sufficiently high accuracy, the functions $K_{1}$ and $\tilde{K}_{2}$ with a relative error of the order less than $10^{-2}$ can be used as the following constants
$K_{1}(\varphi)\approx 1,\qquad\tilde{K}_{2}(\varphi)\approx 1.93\cdot 10^{10}$ (6.5)
It is natural to assume that the initial value $\varphi_{in}$ is significantly greater than $6M_{P}$, that is, at the beginning of evolution, the approximations (6.5) are valid with much greater accuracy than $10^{-2}$.
Figure 4: Plots of $K_{1}(\varphi)$ and $\tilde{K}_{2}(\varphi)$ defined by eqs.(3.9) and (3.10), with the model parameters $\lambda=2.3\cdot 10^{-11}$,   $\xi=\frac{1}{6}$,   $m=0.7GeV$,   $b_{p}\approx 0.5$,   $q^{4}=3\cdot 10^{-10}$ used in Fig.2. In addition the value $b_{k}=1+\mathcal{O}(10^{-5})$ is used.
It is obvious that the structure of the cosmological equations (6.1) and (6.2) differs significantly from the usual ones. Indeed, first of all, in the place where the kinetic energy density should be in the energy density, we find an expression
$K_{eff}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\Big{(}K_{1}-\frac{3}{2}\tilde{K}_{2}\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\Big{)}\cdot\frac{1}{2}\dot{\varphi}^{2},$ (6.6)
which, by analogy with the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$, should be considered as the TMT-effective kinetic energy density. What is very unusual is that $K_{eff}$ and the coefficient before $\ddot{\varphi}$ in eq.(6.2) can be either positive or negative, depending on the values of $\dot{\varphi}^{2}$.
The classical evolution of the Universe implies that from the very beginning the total initial energy density must be positive
$\rho_{in}=K_{eff}^{(in)}+U_{eff}^{(tree)}(\varphi_{in})>0,$ (6.7)
from where we get the upper bound
$\frac{1}{2}\dot{\varphi}_{in}^{2}\lesssim 0.78\cdot U_{eff}^{(tree)}(\varphi_{in})\approx 0.8\cdot 10^{-10}M_{P}^{4}.$ (6.8)
We note that this constraint on the permissible values of $\dot{\varphi}_{in}^{2}$ is somewhat stronger compared to the upper bound imposed by the condition $\zeta>0$, eq.(5.9).
If in classical space-time, at the very beginning of cosmological evolution, the multiplier $\bigl{(}K_{1}-3\tilde{K}_{2}\frac{\dot{\varphi}_{in}^{2}}{2M_{P}^{4}}\bigr{)}$ in front of $\ddot{\varphi}$ in eq.(6.2) is positive, then the classical field $\varphi$ that drives this evolution has regular, "normal" dynamics. But in general, this multiplyer may be negative, that is, the evolution can start not only from the "normal" dynamics, but also from stages with "pathological" dynamics. Another reason for the appearance of "pathological" dynamics may be the case when the TMT-effective kinetic energy density $K_{eff}$ is negative. It is convenient to introduce a notation
$y=\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\lesssim y_{cr}=1.7\cdot 10^{-11},$ (6.9)
and $y_{cr}$ is the critical value of $y$ where the multiplyer $\Bigl{(}K_{1}-3\tilde{K}_{2}\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\Bigr{)}\equiv\Bigl{(}K_{1}-3\tilde{K}_{2}y\Bigr{)}$ in front of $\ddot{\varphi}$ in eq.(6.2) equals zero. Considering the cases when $y\neq y_{cr}$ one can divide eq.(6.2) by the factor standing in front of $\ddot{\varphi}$. This reduces eq.(6.2) to the following form
$\ddot{\varphi}+f_{1}(y)3H\dot{\varphi}+f_{2}(y)\frac{\dot{\varphi}^{2}}{2}+\frac{U_{eff}^{(tree)\prime}(\varphi)}{K_{1}-3\tilde{K}_{2}\cdot y}=0,$ (6.10)
where the notations are introduced
$f_{1}(y)=\frac{K_{1}-\tilde{K}_{2}\cdot y}{K_{1}-3\tilde{K}_{2}\cdot y};\quad f_{2}(y)=\frac{K^{\prime}_{1}-\tilde{K}^{\prime}_{2}\cdot y}{K_{1}-3\tilde{K}_{2}\cdot y}$ (6.11)
The $\varphi$ equation, presented in the form (6.10), allows one to classify, depending on $y$, possible stages of the evolution of the very early Universe according to their types of dynamics.
  • Type I pathological dynamics is realized if $K_{eff}^{(in)}<0$. Then $K_{1}-3\tilde{K}_{2}\frac{\dot{\varphi}_{in}^{2}}{2M_{P}^{4}}<0$ as well. Together with the condition (6.8), ensuring $\rho_{in}>0$, this leads to restrictions on the initial conditions under which the classical evolution of the Universe begins with the first type of pathological dynamics
    $0.35\cdot U_{eff}^{(tree)}(\varphi_{in})<\frac{1}{2}\dot{\varphi}_{in}^{2}<0.8\cdot U_{eff}^{(tree)}(\varphi_{in}).$ (6.12)
  • If $K_{eff}^{(in)}>0$, but $K_{1}-3\tilde{K}_{2}\frac{\dot{\varphi}_{in}^{2}}{2M_{P}^{4}}<0$, then the classical evolution of the Universe begins with type II pathological dynamics, which is realized under the following restrictions on the initial conditions
    $0.17\cdot U_{eff}^{(tree)}(\varphi_{in})\lesssim\frac{1}{2}\dot{\varphi}_{in}^{2}\lesssim 0.35\cdot U_{eff}^{(tree)}(\varphi_{in}).$ (6.13)
  • The condition $K_{1}-3\tilde{K}_{2}\frac{\dot{\varphi}_{in}^{2}}{2M_{P}^{4}}>0$ means that the classical evolution of the Universe can start with normal dynamics only if the following sufficiently strong upper limit on the initial value of $\dot{\varphi}^{2}$ is satisfied
    $\frac{1}{2}\dot{\varphi}_{in}^{2}\lesssim 0.17\cdot U_{eff}^{(tree)}(\varphi_{in})\approx 1.7\cdot 10^{-11}M_{P}^{4},$ (6.14)
    where at the last step we took into account that in the region $\varphi>6M_{P}$ the TMT- effective potential $U_{eff}^{(tree)}(\varphi)$ can be considered flat with a hight that is equal to $U_{eff}^{(tree)}(\varphi_{in})\approx 10^{-10}M_{P}^{4}$ with sufficiently high accuracy. Under the condtion (6.14), the TMT-effective kinetic energy density $K_{eff}^{(in)}$ is positive and is bounded from above by
    $0<K_{eff}^{(in)}\lesssim 8.6\cdot 10^{-2}\cdot U_{eff}^{(tree)}(\varphi_{in})$ (6.15)
    with the upper bound being reached at $\frac{\dot{\varphi}_{in}^{2}}{2M_{P}^{4}}\to 1.7\cdot 10^{-11}$.
Table 1: Types of dynamics at the very beginning of classical cosmological evolution
dynamics
range of sign of sign of sign of active force $w=\frac{p}{\rho}$
$y=\frac{\frac{1}{2}\dot{\varphi}^{2}\cdot 10^{-10}}{U_{eff}^{(tree)}}$ $K_{eff}$ $f_{1}(y)$ $\frac{U_{eff}^{(tree)^{\prime}}}{K_{1}-3\tilde{K}_{2}y}$
pathological I
$0.80.52<y<$ 0.8 $K_{eff}<0$ $f_{1}(y)>0$wrong" $w<-1$
$0.520.35<y<$ 0.52 $K_{eff}<0$ $f_{1}(y)<0$wrong" $w>-1$
pathological II
$0.350.17<y<$ 0.35 $K_{eff}>0$ $f_{1}(y)<0$wrong" $w>-1$
normal
$y<0.17$ $K_{eff}>0$ $f_{1}(y)>0$ regular $w>-1$
The results presented in Table 1 for each type of dynamics (pathological type I, pathological type II and normal dynamics) show: 1) the sign of the TMT-effective kinetic energy density $K_{eff}$; 2) the sign of $f_{1}(y)$ on which the sign of the friction force $f_{1}(y)3H\dot{\varphi}$ depends; 3) the sign of the active force in the $\varphi$ equation (6.10); 4) the sign of $1+w$, where $w$ is the equation of state (EOS) $w=\frac{p}{\rho}$. As can be seen from the table, in fact, pathological dynamics of type I consists of two fundamentally different stages with different EOS. Using the (classical) TMT-effective energy-momentum tensor, eq.(3.11), and the expressions of $K_{1}$ and $\tilde{K}_{2}$ for $\varphi>6M_{P}$, eq.(6.5), we find that
$w=\frac{p}{\rho}=\frac{\tilde{y}-0.97\tilde{y}^{2}-1.07}{\tilde{y}-2.9\tilde{y}^{2}+1.07},\qquad\text{where}\qquad\tilde{y}=y\cdot 10^{10}=\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\cdot 10^{10}.$ (6.16)
The graph of EOS is presented in figure 5, from which it is evident that the evolution of the Universe can begin with $w<-1$. Thus, using the terminology adopted in the models of present-day evolution of the Universe, we come to the conclusion that the classical evolution of the Universe can begin with the stage of phantom dynamics.
Figure 5: Plot of the function (6.16) describing the $\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}$-dependence of the equation of state $w=\frac{p}{\rho}$ in the region $\varphi>6M_{P}$, where the TMT effective potential has a plateau $U_{eff}(\varphi)\approx 10^{-10}M_{P}^{4}$. The condition (6.8), necessary for the total initial energy density to be positive, is manifested in the fact that the line $y=\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\cdot 10^{10}=0.8$ is a vertical asymptote for the function $w(y)$ and $w\to-\infty$ for $y\to 0.8^{-}$. The results presented in eqs.(6.12)-(6.14), together with the shape of the graph $w(y)$, require dividing the interval $0.80<y<$ 0.8 into four regions with different types of dynamics: 1) region $0.80.52<y<$ 0.8 - pathological type I with EoS $w<-1$, i.e. phantom dynamics; 2) region $0.520.35<y<$ 0.52 - pathological type I dynamics with EoS $w>-1$; 3) region $0.350.17<y<$ 0.35 - pathological type II dynamics; 4) region $0.170<y<$ 0.17 - normal dynamics.
At the end of this section, at least brief comments on pathological dynamics are appropriate. It would be interesting to study possible solutions with such initial conditions. Preliminary analysis shows that the pathological stages of evolution are dynamically isolated from the stage with normal dynamics. It may also be noted that such an initial pathological stage of the evolution of the Universe could be responsible for the unknown physics, whose existence has been suggested in the interpretation of the BGV theorem[38]. However, in this paper we will limit ourselves to posing these questions and will study them elsewhere.
6.2 Some important properties of normal dynamics
Considering the normal dynamics, i.e under the condition (6.14), let us turn to the $\varphi$ equation (6.2) represented in the equivalent form (6.10), trying to compare it with the commonly used form of the Klein-Gordon equation in the FLRW Universe.
Using eqs.(6.3), (6.4) and (6.14) it is easy to see that $f_{1}(y)\geq 1$. Therefore, the presence of the factor $f_{1}(y)$ in eq.(6.10) leads to an increase in the usual friction force coefficient $3H$. However, due to the factor $(K_{1}-3\tilde{K}_{2}\cdot y)^{-1}\approx(1-1.93\cdot 10^{10}y)^{-1}$ in the last term of eq.(6.10), an analogous effect occurs, leading to a similar strengthening of the usual active force $-U_{eff}^{(tree)\prime}(\varphi)$.
To estimate the third term in eq.(6.10), we first use the relations (6.3) and (6.4) and find that
$f_{2}(y)=\frac{K^{\prime}_{1}-\tilde{K}^{\prime}_{2}\cdot y}{K_{1}-3\tilde{K}_{2}\cdot y}\approx 185.5\sqrt{\frac{8}{3}}\cdot\frac{f_{1}}{M_{P}}\cdot e^{-\sqrt{\frac{8}{3}}\frac{\varphi}{M_{P}}}.$ (6.17)
For normal dynamics we can use the upper bound (6.14) on $\dot{\varphi}^{2}$, which together with (6.17) and the condition $\varphi>6M_{P}$ gives an estimate for the third term in eq.(6.10)
$f_{2}\frac{\dot{\varphi}^{2}}{2}\approx 307.8\frac{f_{1}}{M_{P}}\cdot e^{-\sqrt{\frac{8}{3}}\frac{\varphi}{M_{P}}}\sqrt{0.17\cdot U_{eff}^{(tree)}}\frac{|\dot{\varphi}|}{\sqrt{2}}<5\cdot 10^{-3}\cdot\frac{f_{1}}{M_{P}}\sqrt{U_{eff}^{(tree)}}\cdot|\dot{\varphi}|.$ (6.18)
Since $K_{eff}>0$, it follows from eq.(6.1) that
$H=\frac{1}{\sqrt{3}M_{P}}\sqrt{K_{eff}+U_{eff}^{(tree)}}>\frac{1}{\sqrt{3}M_{P}}\sqrt{U_{eff}^{(tree)}},$ (6.19)
Then for the contribution of the second term in eq.(6.10) we obtain
$|3f_{1}H\dot{\varphi}|>\sqrt{3}\frac{f_{1}}{M_{P}}\sqrt{U_{eff}^{(tree)}}|\dot{\varphi}|.$ (6.20)
Thus, the contribution of $f_{2}\frac{\dot{\varphi}^{2}}{2}$ in eq.(6.10) is at least $2.9\cdot 10^{-3}$ times smaller than $|3f_{1}H\dot{\varphi}|$, and therefore it will be omitted in the further study of normal dynamics.
Applying the result (6.15) to the Friedmann equation (6.1), we see that it can be viewed as a modification of the condition (1.1) required for inflation to begin (it is assumed that $K_{eff}^{(in)}$ is greater than the initial TMT-effective gradient energy density). Thus,, we arrive at the following two fundamentally new conclusions:
  • classical cosmological evolution with normal dynamics can begin only if
    $\frac{1}{2}\dot{\varphi}_{in}^{2}<0.17\cdot U_{eff}^{(tree)}(\varphi_{in})$;
  • classical cosmological evolution with normal dynamics
    begins with the inflationary regime.
    Thus, in the studied Two-Measure Standard Model of Higgs inflation, the problem of initial conditions for inflation does not arise.
6.3 About the slow-roll phase of inflation,
taking into account the K-essence type structure.
Now, using the results of the previous subsection, we are going to show that the slow-roll approximation is valid during inflation. Furthermore, we want to see what changes occur compared to what we had in section 3.2, where the K-essence structure was completely ignored. Considering the shape of $K_{1}(\varphi)$ and $K_{2}(\varphi)$, one can expect that this effect on inflation at $\varphi>6M_{P}$ is minor. But at the stage when inflation is approaching its end, which happens at $\varphi<6M_{P}$ , where $K_{1}(\varphi)$ and $K_{2}(\varphi)$ noticeably decrease, this influence can be significant.
As we already know, the initial conditions (6.14) and (6.15) not only ensure the onset of cosmological evolution with normal dynamics, but also the evolution starts with inflation. Here it is convenient to return from eq.(6.10) to its original form (6.2). We explore the region $\varphi>6M_{P}$, where, as we also know, the contribution of the third term in eq. (6.2) is negligible compared to the second term. From eq.(6.1) it follows that the first slow-roll condition requires that
$K_{eff}=K_{1}\cdot\frac{1}{2}\dot{\varphi}^{2}-\frac{3}{2}\cdot\tilde{K}_{2}\frac{\dot{\varphi}^{4}}{4M_{P}^{4}}\ll U_{eff}^{(tree)}(\varphi)$ (6.21)
At the very beginning of inflation, due to the frictional force $\propto 3H\dot{\varphi}$, $\dot{\varphi}^{2}$ decreases, and $\dot{\varphi}^{4}$ decreases even faster. Therefore, not only does $K_{eff}$ remain positive, but the term $\frac{3}{2}\cdot\tilde{K}_{2}\frac{\dot{\varphi}^{4}}{4M_{P}^{4}}$ quickly becomes subdominant compared to $K_{1}\cdot\frac{1}{2}\dot{\varphi}^{2}$. Hence, the first slow-roll condition very quickly reduces to the standard one
$K_{1}\cdot\frac{1}{2}\dot{\varphi}^{2}\approx\frac{1}{2}\dot{\varphi}^{2}\ll U_{eff}^{(tree)}(\varphi)\approx 10^{-10}M_{P}^{4},$ (6.22)
which agrees with the condition (6.14) under which normal dynamics, and therefore inflation, exists.
Thanks to the known shapes of $K_{1}(\varphi)$ and $\tilde{K}_{2}(\varphi)$ when $\varphi>6M_{P}$ and using (6.22) we get
$\tilde{K}_{2}\cdot\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\ll\frac{\tilde{K}_{2}}{K_{1}}\frac{U_{eff}^{(tree)}(\varphi)}{M_{P}^{4}}\approx 2.$ (6.23)
Figure 6: Graphs of the functions $\epsilon(\varphi)$ and $\epsilon_{k}(\varphi)$ defined in eqs.(3.21) and (6.26), respectively.
Therefore, in the $\varphi$-equation (6.2) one can neglect the contributions $\propto\tilde{K}_{2}(\varphi)\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}$ compared to $K_{1}(\varphi)\approx 1$. Hence, the second slow-roll condition reduces to $|K_{1}(\varphi)\ddot{\varphi}|\ll|3HK_{1}(\varphi)\dot{\varphi}|$, that for $\varphi>6M_{P}$ coincides with the usual one $|\ddot{\varphi}|\ll|3H\dot{\varphi}|$. Thus, the cosmological equations (6.1) and (6.2) in the slow-roll approximation read
$H^{2}=\frac{1}{3M_{P}^{2}}U_{eff}^{(tree)}(\varphi),$ (6.24)
$3HK_{1}(\varphi)\dot{\varphi}=-U_{eff}^{(tree)\prime}(\varphi).$ (6.25)
One can show [^15] It is interesting that a somewhat similar equations appear in the Palatini-Higgs inflation considered in the model with non-minimal derivative coupling [57]. that in this slow-roll approximation the inflationary parameters $\epsilon$ and $\eta$, eq.(3.21), and the corresponding flatness conditions are replaced by
$\epsilon_{k}=\frac{M_{P}^{2}}{2K_{1}}\Biggl{(}\frac{U_{eff}^{(tree)\prime}}{U_{eff}^{(tree)}}\Biggr{)}^{2}\ll 1,\qquad|\eta_{k}|=\frac{M_{P}^{2}}{U_{eff}^{(tree)}\sqrt{K_{1}}}\Biggl{|}\Biggl{(}\frac{U_{eff}^{(tree)\prime}}{\sqrt{K_{1}}}\Biggr{)}^{\prime}\Biggr{|}\ll 1,$ (6.26)
where we used the subscript $k$ to indicate that the TMT-effective K-essence-type structure was used to derive these parameter expressions. Since the deviation of $K_{1}(\varphi)$ from unity becomes significant only for $\varphi<6M_{P}$, the difference between $\epsilon$ and $\epsilon_{k}$ can only appear at the last stage of inflation. To illustrate the expected effect, figure 6 shows graphs of the $\varphi$-dependencies of $\epsilon$ and $\epsilon_{k}$. As we can see, $\epsilon_{k}$ approaches unity at higher $\varphi$. Thus, one of the expected effects of the K-essence type structure is an earlier end of inflation compared to the simplified model of section 3.
7 Discussion
In the discussion of the basic ideas and main results of the paper, it is worth concentrating on the role played by the combination of the presence of two volume measures in the TMSM with the flatness of the TMT-effective potential.
To understand how Higgs inflation becomes possible at small $\xi$, we should start by recalling the special properties of the model that allow us to overcome the apparent contradiction between physics at inflationary-scale energies and the energies of accelerator experiments. The choice of a small value of the non-minimal coupling constant $\xi$ can be consistent with the CMB data only if the primordial model parameter $\lambda$ of the Higgs field selfcoupling is very small: as was shown, $\lambda=2.3\cdot 10^{-11}$ if $\xi=1/6$. But such a value of $\lambda$ is completely unsuitable for the implementation of the GWS theory. The resolution of this contradiction in the studied TMSM turns out to be possible only due to two volume measures in the primordial action. The main consequence of this is that $\lambda$ becomes a kind of running classical coupling constant, the value of which at energies near the Higgs field vacuum turns out to be approximately $10^{10}$ times larger than during the inflation stage. The very possibility of such a huge difference is the result of the fact that the value of the coefficient $b_{k}$ in the volume element $(b_{k}\sqrt{-g}-\Upsilon)d^{4}x$ is chosen extremely close to 1: $b_{k}=1+\mathcal{O}(10^{-5})$. The coefficient $b_{k}$ can be interpreted as a factor describing the relative contributions of the volume measure densities $\sqrt{-g}$ and $\Upsilon$ to the total volume element. Therefore, such a small deviation of $b_{k}$ from 1 may indicate the existence of a more fundamental theory in which the original balance between $\sqrt{-g}$ and $\Upsilon$ is broken by some weak effect. Thus, in TMSM the contradiction described above is resolved in a fairly natural way.
Now let us return to the problem mentioned in the introductory section 1 that the initial conditions for inflation are unnatural. Such claims are widespread in the literature and are usually based on statistical estimates that the initial conditions (1.1) required for inflation to begin are exponentially rare. This argument became even more convincing when it turned out that models with plateau-like potentials are the most preferable in light of the Planck’s CMB results. As the authors of the paper [30] rightly noted, for a flat potential of height $\sim 10^{-10}M_{P}^{4}$ one would expect that initial values of the kinetic and gradient energy densities exceeding $\sim 10^{-10}M_{P}^{4}$, up to the Planck scale, are the most probable; and conversely, values of these energies satisfying the conditions (1.1) and necessary for the onset of inflation have a much lower probability. However, it is easy to see that in the context of TMSM this argument turns out to be completely inapplicable. First of all, given the K-essence type structure, we know that according to condition (6.8), the initial energy density $\rho_{in}$ turns out to be negative if $\frac{1}{2}\dot{\varphi}_{in}^{2}$ exceeds $0.8\cdot 10^{-10}M_{P}^{4}$. But one can also make a more fundamental argument that is in no way connected with the K-essence type structure. To do this, we recall that, as discussed in detail in section 5, in our Universe the ratio of two volume measures described by the scalar function $\zeta$ must be strictly positive. In section 5.2, analyzing the constraint (5.2) determining the values of $\zeta$, we found that for the beginning of the classical evolution of the Universe it is necessary that the initial values $\dot{\varphi}_{in}$ and $\nabla\varphi_{in}$ satisfy the upper bounds (5.9) and (5.10), respectively. From this analysis it follows that $\zeta$ becomes negative if we wish to choose initial values $\dot{\varphi}_{in}$ and $\nabla\varphi_{in}$ that exceed these upper bounds. And this, firstly, contradicts the conditions of applicability of the equations obtained from the principle of least action and, secondly, means changing the orientation of the spacetime manifold of our Universe to the opposite.
The absence of the problem of initial conditions for Higgs inflation in the developed model is part of a more general result, namely, that inflation can be preceded by other, rather exotic initial stages of classical cosmological evolution. Since this result has fundamental theoretical significance, it is worth discussing its origin on a qualitative level. Since at $\varphi>6M_{P}$ the TMT-effective potential has a plateau of almost constant height, the possibility of the onset of cosmological evolution turns out to be independent of the initial value $\varphi_{in}>6M_{P}$. Therefore, from the pair of initial conditions $(\varphi_{in},\dot{\varphi}_{in})$ the possibility of the onset of evolution turns out to depend only on $\dot{\varphi}_{in}$. On the other hand, the TMT primordial action (2.1) leads to a TMT-effective action (3.4)-(3.10) having a K-essence-type structure: the kinetic part of the TMT-effective Lagrangian (3.5) is a quadratic form in $\dot{\varphi}^{2}$. This structure implies that the condition of positive initial total energy density, $\rho_{in}>0$, allows the initial TMT-effective kinetic energy density to be negative (although $\dot{\varphi}^{2}>0$). This makes it possible that the evolution of the Universe starts with an equation of state $w=\frac{p}{\rho}<-1$, i.e. proceeds according to the dynamics known in dark energy models as phantom. Other pathological initial states with $w>-1$ are also possible. Cosmological solutions describing the stages with pathological dynamics are of great interest and will be studied elsewhere. However, if the evolution starts with normal dynamics, then from the very beginning it proceeds in the inflationary regime.